### v2

# Univariate-Guided Interaction Modeling

Abstract
We propose a procedure for sparse regression with pairwise interactions, by generalizing the Univariate Guided Sparse Regression (UniLasso) methodology. A central contribution is our introduction of a concept of univariate (or marginal) interactions. Using this concept, we propose two algorithms— uniPairs and uniPairs-2stage—, and evaluate their performance against established methods, including Glinternet and Sprinter. We show that our framework yields sparser models with more interpretable interactions. We also prove support recovery results for our proposal under suitable conditions.
1 Introduction
We consider the problem of modeling pairwise interactions between features where the target $Y\in\mathbb{R}^{n}$ follows the model
$Y=\beta^{*}_{0}+X^{T}\beta^{*}+Z^{T}\gamma^{*}+\epsilon$
with $X\in\mathbb{R}^{n\times p}$ denoting the design matrix of main effects, $Z$ encoding the interactions derived from $X$ and $\epsilon$ being random noise. The goal is to identify a sparse subset of both main effects and interactions that has strong predictive power with respect to $Y$, particularly in the high-dimensional regime $p\gg n$.
This work is based on the Univariate Guided Sparse Regression (UniLasso) framework introduced in Chatterjee et al. [3], and extends the idea of univariate guidance to both main effects and interaction terms. Our primary objectives are to:
  • achieve competitive prediction error and sparser models than existing methods such as Sprinter, HierNet, and Glinternet.
  • deliver pairwise interactions which are more interpretable than those produced by competing procedures.
  • focus on the the high-dimensional regime where $p\gg n$, and design an algorithm whose time and space complexity is sub-quadratic in $p$, enabling scalability to high-dimensional settings, or which can be parallelized on modern hardware to avoid the quadratic cost.
More broadly, the research problem is a subclass of the following: given a set of original features (main effects), and a method for engineering new ones (e.g. pairwise interactions), how can we efficiently select a subset of these engineered features so that when combined with the original chosen features, predictive performance improves? This is especially important when the number of engineered features far exceeds the number of original ones, as in the case of pairwise interactions.
Interaction modeling has diverse applications. In genetics, interactions between genes (epistasis) can reveal mechanisms responsible for complex traits. In medicine, the simultaneous presence of two symptoms may enhance (positive interaction) or cancel (negative interaction) diagnostic information. In recommender systems, user-item interactions are important for personalization.
In such settings, methods like the All Pairs Lasso (APL) [^1] This method adds all pairwise products to a linear model, and fits the model using the lasso. become computationally expensive and tend to favor spurious interaction terms over true main effects. APL requires $O(np^{2})$ memory and performs multiple passes through coordinate descent on $O(p^{2})$ interaction terms, making it not scalable.
Several methodological paradigms have been proposed for modeling interactions. These include multi-stage procedures such as Sprinter and Regularized regression frameworks like HierNet[2] and Glinternet[7]. Our proposal is closest in spirit to Sprinter[12]. This method regresses out main effects, then scans all candidate interaction terms for their correlation with the resulting residuals, and finally regresses the residuals on the selected interactions.
What sets our proposal apart is our definition and use of marginal pairwise interactions. We use these marginal interactions to screen the large number of candidate pairwise products in building our model. This improves the accuracy of the true model recovery and delivers interactions that are more credible and interpretable.
This paper is organized as follows. Section 2 introduces our Univariate-guided procedures for interaction modeling and describes the high level ideas behind uniPairs and uniPairs-2stage. Section 2.3 presents a motivating application to HIV-mutation data. Section 3 gives the full algorithmic details, including the TripletScan screening step and the UniLasso/Lasso fits. Section 4 summarizes the main ideas from related work on interaction modeling including Sprinter, Glinternet, HierNet and Group Lasso approaches. Section 5 shows more findings on the HIV-mutation data. Section 6 reports the results of a simulation study comparing uniPairs and uniPairs-2stage to the existing methods Sprinter and Glinternet. Section 7.1 states our theoretical results about support recovery and $\ell_{\infty}$ estimation error of the coefficients in the UniLasso step for both methods, extending the UniLasso analysis in Chatterjee et al. [3], with proofs deferred to Appendix Appendix A: Proofs. Section 7.2 explains the statistical motivation behind our screening rule, its connection with conditional sure independence screening and the rational for the largest log-gap thresholding rule. We conclude in Section 8 with a brief summary and a discussion of possible future work. Appendix Appendix B: Generalization to Binomial GLM and Cox survival model shows how to extend uniPairs and uniPairs-2stage to the Binomial generalized linear model and the Cox proportional hazards model.
2 Our Proposed Algorithms
2.1 Setup and Notation
Let $X\in\mathbb{R}^{n\times p}$ be the design matrix and $Y\in\mathbb{R}^{n}$ be the response vector. For $j\in[p]$, let $X_{j}$ denote the $j$th column of $X$. For $j<k$, define the interaction column $X_{j}\odot X_{k}\in\mathbb{R}^{n}$. Let
$\mathcal{P}=\{(j,k)\in[p]^{2}\ \,|\,\ j<k\}\quad\text{and}\quad Z=\big(X_{j}\odot X_{k}\big)_{(j,k)\in\mathcal{P}}\in\mathbb{R}^{n\times\binom{p}{2}}$
2.2 The high level idea
The steps of our uniPairs procedure are as follows:
(a) For each $j,k$, fit a least squares model of $Y$ on the triplet $(X_{j},X_{k},X_{j}\odot X_{k})$ and measure the contribution to the fit due to $X_{j}\odot X_{k}$. Retain the pairs that have contributions greater than some data-adaptive threshold using the largest log-gap rule. Note that the threshold is not an added hyperparameter, but is is data-adaptive.
(b) Apply the uniLasso algorithm with target $Y$ to all individual features and the pairs that pass the screen in step (a).
We think of the feature pairs that pass the screen in (a) as displaying marginal interaction. If they are chosen for the final model in step (b), these interactions are more credible and interpretable than those pairs with weak marginal interaction.
Our uniPairs-2stage procedure is very similar, except that we fit main effects as a first stage:
(a) Apply UniLasso to the individual features and compute the residual $R$.
(b) For each $j,k$, fit a least squares model of $Y$ the triplet $(X_{j},X_{k},X_{j}\odot X_{k})$ and measure the contribution to the fit due to $X_{j}\odot X_{k}$.
(c) Apply the Lasso with target $R$ to the pairs that pass the screen in step (b).
The final model is the sum of the two models obtained from steps (a) and (c).
Notice that the triplet regression in both procedures use $Y$ as their target; perhaps surprisingly in the uniPairs-2stage procedure where the residual $R$ might seem a more appropriate target. The reason is that we want to find interaction pairs for our multivariate model that also display marginal interactions with $Y$.
We call our procedure “Univariate guided” because the main effects estimation in Step (a) uses univariate guidance, and the contribution of the $j,k$ interaction is measured by the linear covariance between the response and the interaction conditional on its two main effects (section 7.1).
Before giving details of our proposed method, we show a motivating example.
2.3 Example: HIV mutation data
As an example, [10] studied six nucleoside reverse transcriptase inhibitors that are used to treat HIV-1. The target of these drugs can become resistant through mutation, and they compared a collection of models for predicting the log susceptibility, a measure of drug resistance based on the location of mutations. We chose one of the inhibitors, with a total of $n=1005$ samples and $p=211$ mutation sites. We retained features with $\geq 5\%$ ones, leaving $69$ features. uniPairs-2stage and uniPairs chose 24 and 23 main effects respectively, and four interaction pairs. The marginal interactions for the four chosen pairs are displayed in Figure 1. All four pairs show strong interaction effects.
Figure 1: Interactions found in the HIV data. The vertical axis shows the average value of the interactions, as measured by $\widehat{\beta}_{jk,jk}$ from the pairwise model (1)
3 Details of our proposed algorithms
3.1 Summary
Algorithm 1 TripletScan
0: Standardized design matrix $X\in\mathbb{R}^{n\times p}$, response $Y\in\mathbb{R}^{n}$, pair index set $\mathcal{P}$.
1:for each $(j,k)\in\mathcal{P}$ do
2:  Fit local OLS: $Y=\beta_{0,jk}+\beta_{j,jk}X_{j}+\beta_{k,jk}X_{k}+\beta_{jk,jk}X_{j}\odot X_{k}+\varepsilon$.
3:  Record the two-sided t-test $p$-value $p_{jk}$ for $\beta_{jk,jk}$.
4:end for
5: Sort $\{p_{jk}\}$ increasingly, set $\ell_{r}=\log\widehat{p}_{(r)}$, and apply the largest log-gap rule:
$\widehat{r}=\arg\max_{1\leq r<M}(\ell_{r+1}-\ell_{r}),\quad\widehat{\Gamma}=\{(j,k)\in\mathcal{P}:p_{jk}\leq p_{(\widehat{r})}\}$.
5: Selected interactions $\widehat{\Gamma}$.
Algorithm 2 uniPairs-2stage
0: Design matrix $X\in\mathbb{R}^{n\times p}$, response $Y\in\mathbb{R}^{n}$, hierarchy level $h\in\{\text{strong},\text{weak},\text{none}\}$.
1: Standardize each column of $X$.
2: Fit UniLasso on $(X,Y)$ to obtain main-effects active set $\widehat{S}_{M}$ and prevalidated predictions $\widehat{Y}^{(1)}_{\mathrm{PV}}$.
3: Run TripletScan on $(X,Y)$.
4: Restrict eligible pairs $\mathcal{E}$ based on hierarchy level $h$ and $\widehat{S}_{M}$
5: Compute residual $R=Y-\widehat{Y}^{(1)}_{\mathrm{PV}}$.
6: Fit a Lasso of $R$ on the selected interactions $\{X_{j}\odot X_{k}:(j,k)\in\widehat{\Gamma}\cap\mathcal{E}\}$.
7: Recover coefficients on the original scale and get active sets $\widehat{S}_{M}^{\textit{final}}$ and $\widehat{S}_{I}^{\textit{final}}$.
7: Predictive function $\widehat{f}(x)=\widehat{\alpha}_{0}+\sum_{j\in\widehat{S}_{M}^{\textit{final}}}\widehat{\alpha}_{j}x_{j}+\sum_{(j,k)\in\widehat{S}_{I}^{\textit{final}}}\widehat{\alpha}_{jk}x_{j}x_{k}.$
Algorithm 3 uniPairs
0: Design matrix $X\in\mathbb{R}^{n\times p}$, response $Y\in\mathbb{R}^{n}$.
1: Standardize each column of $X$.
2: Run TripletScan on $(X,Y)$ to obtain interaction set $\widehat{\Gamma}$.
3: Form augmented design $\widetilde{X}=[X,X_{\widehat{\Gamma}}$], where $X_{\widehat{\Gamma}}=\{X_{j}\odot X_{k}:(j,k)\in\widehat{\Gamma}\}$.
4: Fit UniLasso on $(\widetilde{X},Y)$.
5: Recover coefficients on the original scale and get active sets $\widehat{S}_{M}$ and $\widehat{S}_{I}$.
5: Predictive function $\widehat{f}(x)=\widehat{\alpha}_{0}+\sum_{j\in\widehat{S}_{M}}\widehat{\alpha}_{j}x_{j}+\sum_{(j,k)\in\widehat{S}_{I}}\widehat{\alpha}_{jk}x_{j}x_{k}.$
In practice, we suggest uniPairs-2stage as a default when main-effects are believed to be present and strong/weak hierarchy holds. uniPairs can be seen as a flexible alternative when departures from hierarchy are expected.
3.2 The uniPairs procedure
For each $j\in[p]$, let $\mu_{j}=\frac{1}{n}\sum_{i=1}^{n}X_{ij}$ and $\sigma_{j}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{ij}-\mu_{j})^{2}$. Define
$\widetilde{X}_{ij}=\frac{X_{ij}-\mu_{j}}{\sigma_{j}}$
For each $(j,k)\in\mathcal{P}$, fit the OLS model
$Y=\beta_{0,jk}+\beta_{j,jk}\widetilde{X}_{j}+\beta_{k,jk}\widetilde{X}_{k}+\beta_{jk,jk}\widetilde{X}_{j}\odot\widetilde{X}_{k}+\varepsilon$ (1)
Record $\widehat{\beta}_{jk}=\widehat{\beta}_{jk,jk}$ and
$\widehat{p}_{jk}=\text{two-sided $t$-test $p$-value for }\widehat{\beta}_{jk}$
As opposed to uniPairs-2stage, we can only take the eligible set $\mathcal{E}=\mathcal{P}$.
Sort $\{\widehat{p}_{jk}:(j,k)\in\mathcal{E}\}$ increasingly as
$\widehat{p}_{(1)}\leq\cdots\leq\widehat{p}_{(M)}$
with $M=|\mathcal{E}|$. Let $\widehat{p}^{\circ}_{(r)}=\max\{\widehat{p}_{(r)},10^{-20}\}$ and set $\ell_{r}=\log\widehat{p}^{\circ}_{(r)}$. Choose
$\widehat{r}\ \in\ \arg\max_{1\leq r<M}\ (\ell_{r+1}-\ell_{r}),\qquad\widehat{\Gamma}\ =\ \big\{(j,k)\in\mathcal{E}:\ \widehat{p}_{jk}\leq\widehat{p}_{(\widehat{r})}\big\}$
For each $j\in[p]$, fit the Univariate OLS model $Y=\beta_{0,j}^{\mathrm{uni}}+\beta_{1,j}^{\mathrm{uni}}\widetilde{X}_{j}+\epsilon.$
For each $(j,k)\in\widehat{\Gamma}$, fit the Univariate OLS model $Y=\beta_{0,jk}^{\mathrm{uni}}+\beta_{1,jk}^{\mathrm{uni}}\widetilde{X}_{j}\odot\widetilde{X}_{k}+\epsilon.$ For each $i\in[n]$, compute the leave-one-out predictions
$\widehat{\eta}_{j}^{(-i)}=\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}+\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}\widetilde{X}_{ij},\qquad\widehat{\eta}_{jk}^{(-i)}=\widehat{\beta}_{0,jk}^{(-i)\mathrm{uni}}+\widehat{\beta}_{1,jk}^{(-i)\mathrm{uni}}\widetilde{X}_{ij}\widetilde{X}_{ik}$
Given a penalty level $\lambda>0$, solve
[…]
$\forall(j,k)\in\widehat{\Gamma}\quad\theta_{jk}^{s}$ $\geq 0$
Select $\lambda$ by $K$-fold cross-validation and refit at the chosen value. Denote a solution by $(\widehat{\theta}_{0}^{s},\widehat{\theta}^{s})$ and define
$\widehat{\beta}_{j}^{s}\ =\ \widehat{\theta}_{j}^{s}\widehat{\beta}_{1,j}^{\mathrm{uni}},\qquad\widehat{\beta}_{jk}^{s}\ =\ \widehat{\theta}_{jk}^{s}\widehat{\beta}_{1,jk}^{\mathrm{uni}},\qquad\widehat{\beta}_{0}^{s}\ =\ \widehat{\theta}_{0}^{s}+\sum_{j=1}^{p}\widehat{\theta}_{j}^{s}\widehat{\beta}_{0,j}^{\mathrm{uni}}+\sum_{(j,k)\in\widehat{\Gamma}}\widehat{\theta}_{jk}^{s}\widehat{\beta}_{0,jk}^{\mathrm{uni}}$
Convert back to the original scale
$\widehat{\beta}_{jk}=\frac{\widehat{\beta}_{jk}^{(s)}}{\sigma_{j}\sigma_{k}},\qquad\widehat{\beta}_{j}=\frac{\widehat{\beta}_{j}^{(s)}}{\sigma_{j}}-\frac{1}{\sigma_{j}}\sum_{k\neq j}\frac{\widehat{\beta}_{jk}^{(s)}}{\sigma_{k}}\,\mu_{k},$
$\widehat{\beta}_{0}=\widehat{\beta}_{0}^{(s)}-\sum_{j=1}^{p}\frac{\widehat{\beta}_{j}^{(s)}\mu_{j}}{\sigma_{j}}+\sum_{(j,k)\in\widehat{\Gamma}}\frac{\widehat{\beta}_{jk}^{(s)}\mu_{j}\mu_{k}}{\sigma_{j}\sigma_{k}}$
Define the active sets as
$\widehat{S}_{M}=\{j\in[p]:\ \widehat{\beta}_{j}\neq 0\},\qquad\widehat{S}_{I}=\{(j,k)\in\widehat{\Gamma}:\ \widehat{\beta}_{jk}\neq 0\}$
The fitted model is
$\widehat{f}(x)\;=\;\widehat{\beta}_{0}+\sum_{j\in\widehat{S}_{M}}\widehat{\beta}_{j}\,x_{j}+\sum_{(j,k)\in\widehat{S}_{I}}\widehat{\beta}_{jk}\,x_{j}x_{k}$
3.3 The uniPairs-2stage procedure
For each $j\in[p]$, let $\mu_{j}=\frac{1}{n}\sum_{i=1}^{n}X_{ij}$ and $\sigma_{j}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{ij}-\mu_{j})^{2}$. Define
$\widetilde{X}_{ij}=\frac{X_{ij}-\mu_{j}}{\sigma_{j}}$
For each $j\in[p]$, fit the Univariate OLS model $Y=\beta_{0,j}^{\mathrm{uni}}+\beta_{1,j}^{\mathrm{uni}}\widetilde{X}_{j}+\epsilon.$
For each $i\in[n]$, compute the leave-one-out predictions
$\widehat{\eta}_{j}^{(-i)}=\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}+\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}\widetilde{X}_{ij}$
Given a penalty level $\lambda_{1}>0$, solve
[…]
Select $\lambda_{1}$ by $K$-fold cross-validation, get the prevalidated predictions $\widehat{Y}^{(1)}_{\mathrm{PV}}\in\mathbb{R}^{n}$ at the chosen $\lambda_{1}$ and refit on the full data at this value.
Denote a solution by $(\widehat{\theta}_{0}^{s},\widehat{\theta}^{s})$ and define
$\widehat{\beta}_{j}^{s}\ =\ \widehat{\theta}_{j}^{s}\widehat{\beta}_{1,j}^{\mathrm{uni}},\qquad\widehat{\beta}_{0}^{s}\ =\ \widehat{\theta}_{0}^{s}+\sum_{j=1}^{p}\widehat{\theta}_{j}^{s}\widehat{\beta}_{0,j}^{\mathrm{uni}}$
Let $\widehat{S}_{M}^{(1)}=\{j\in[p]:\ \widehat{\beta}_{j}^{s}\neq 0\}$. For each $(j,k)\in\mathcal{P}$, fit the OLS model
$Y=\beta_{0,jk}+\beta_{j,jk}\widetilde{X}_{j}+\beta_{k,jk}\widetilde{X}_{k}+\beta_{jk,jk}\widetilde{X}_{j}\odot\widetilde{X}_{k}+\varepsilon$
Record $\widehat{\beta}_{jk}=\widehat{\beta}_{jk,jk}$ and
$\widehat{p}_{jk}=\text{two-sided $t$-test $p$-value for }\widehat{\beta}_{jk}$
Given a hierarchy regime $h\in\{\text{strong},\text{weak},\text{none}\}$, define the eligible set
$\mathcal{E}=\begin{cases}\{(j,k)\in\mathcal{P}:\ j\in\widehat{S}_{M}^{(1)}\ \text{and}\ k\in\widehat{S}_{M}^{(1)}\}&\text{ if }h=\text{strong},\\ \{(j,k)\in\mathcal{P}:\ j\in\widehat{S}_{M}^{(1)}\ \text{or}\ k\in\widehat{S}_{M}^{(1)}\}&\text{ if }h=\text{weak},\\ \mathcal{P}&\text{ if }h=\text{none}.\end{cases}$
If $\widehat{S}_{M}^{(1)}=\emptyset$, take $h=\text{none}$. Sort $\{\widehat{p}_{jk}:(j,k)\in\mathcal{E}\}$ increasingly as
$\widehat{p}_{(1)}\leq\cdots\leq\widehat{p}_{(M)}$
with $M=|\mathcal{E}|$. Let $\widehat{p}^{\circ}_{(r)}=\max\{\widehat{p}_{(r)},10^{-20}\}$ and set $\ell_{r}=\log\widehat{p}^{\circ}_{(r)}$. Choose
$\widehat{r}\ \in\ \arg\max_{1\leq r<M}\ (\ell_{r+1}-\ell_{r}),\qquad\widehat{\Gamma}\ =\ \big\{(j,k)\in\mathcal{E}:\ \widehat{p}_{jk}\leq\widehat{p}_{(\widehat{r})}\big\}$
Let the prevalidated residual be
[…]
Given a penalty level $\lambda_{2}>0$, solve
[…]
Select $\lambda_{2}$ by $K$-fold cross-validation, then convert back to the original scale
$\widehat{\beta}_{jk}=\frac{\widehat{\alpha}_{jk}^{(s)}}{\sigma_{j}\sigma_{k}},\qquad\widehat{\beta}_{j}=\frac{\widehat{\beta}_{j}^{(s)}}{\sigma_{j}}-\frac{1}{\sigma_{j}}\sum_{k\neq j}\frac{\widehat{\alpha}_{jk}^{(s)}}{\sigma_{k}}\,\mu_{k},$
$\widehat{\beta}_{0}=\widehat{\beta}_{0}^{(s)}+\widehat{\alpha}_{0}^{(s)}-\sum_{j=1}^{p}\frac{\widehat{\beta}_{j}^{(s)}\mu_{j}}{\sigma_{j}}+\sum_{(j,k)\in\widehat{\Gamma}}\frac{\widehat{\alpha}_{jk}^{(s)}\mu_{j}\mu_{k}}{\sigma_{j}\sigma_{k}}$
Define the active sets as
$\widehat{S}_{M}=\{j\in[p]:\ \widehat{\beta}_{j}\neq 0\},\qquad\widehat{S}_{I}=\{(j,k)\in\widehat{\Gamma}:\ \widehat{\beta}_{jk}\neq 0\}$
The fitted model is
$\widehat{f}(x)\;=\;\widehat{\beta}_{0}+\sum_{j\in\widehat{S}_{M}}\widehat{\beta}_{j}\,x_{j}+\sum_{(j,k)\in\widehat{S}_{I}}\widehat{\beta}_{jk}\,x_{j}x_{k}$
3.4 Effect of Standardizing the Main Effects
Consider again the true model with main and pairwise interaction terms:
$Y_{i}=\beta_{0}^{*}+\sum_{j=1}^{p}\beta_{j}^{*}\,X_{ij}+\sum_{1\leq j<k\leq p}\gamma_{jk}^{*}\,X_{ij}X_{ik}+\varepsilon_{i}$
Define the standardized covariates
$X_{ij}^{s}=\frac{X_{ij}-\mathbb{E}[X_{ij}]}{\sqrt{\mathrm{Var}(X_{ij})}}=\frac{X_{ij}-\mu_{j}}{\sigma_{j}}$
and rewrite the model in terms of standardized variables as
$Y=\beta_{0}^{*,s}+\sum_{j=1}^{p}\beta_{j}^{*,s}\,X_{j}^{s}+\sum_{1\leq j<k\leq p}\gamma_{jk}^{*,s}\,X_{j}^{s}X_{k}^{s}$
Then
$\beta_{0}^{*}=\beta_{0}^{*,s}-\sum_{j=1}^{p}\frac{\beta_{j}^{*,s}\mu_{j}}{\sigma_{j}}+\sum_{1\leq j<k\leq p}\frac{\gamma_{jk}^{*,s}}{\sigma_{j}\sigma_{k}}\,\mu_{j}\,\mu_{k}$
and for each $j\in[p]$
$\beta_{j}^{*}=\frac{\beta_{j}^{*,s}}{\sigma_{j}}-\frac{1}{\sigma_{j}}\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{p}\frac{\gamma_{jk}^{*,s}}{\sigma_{k}}\,\mu_{k}$ (2)
and
$\gamma_{jk}^{*}=\frac{\gamma_{jk}^{*,s}}{\sigma_{j}\sigma_{k}}\qquad\forall\,1\leq j<k\leq p$
By convention,
$\gamma_{kj}^{*}=\gamma_{jk}^{*},\qquad\gamma_{jj}^{*}=0$
During the fitting procedure, the main effects $X_{j}$ are standardized to $X_{j}^{s}$ and the uniPairs-2stage / uniPairs model is fit on the $X_{j}^{s}$’s.
When the fitting is complete, the fitted linear model in terms of the standardized covariates $X_{j}^{s}$ is rewritten back in terms of the original covariates $X_{j}$. Hence, there are two sources of hierarchy:
  1. 1.
    The explicit enforcement of hierarchy during fitting — this acts on the $Z_{j}$’s. This holds when the option hierarchy in uniPairs-2stage is set to strong or weak. By default, it is set to None, in which case no hierarchy is enforced and the triplet scans are performed on all $\binom{p}{2}$ pairs of $X_{j}^{s}$’s.
  2. 2.
    The implicit hierarchy induced by re-expressing the fitted model in the original variables $X_{j}$. This comes from the term $\frac{1}{\sigma_{j}}\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{p}\frac{\gamma_{jk}^{*,s}}{\sigma_{k}}\,\mu_{k}$ in 2.
Therefore, unless the $X_{j}$’s are already mean-zero, this final conversion step will (almost surely) enforce hierarchy automatically.
4 Related work
Here, we summarize the main ideas from related work on interaction modeling.
4.1 On the need for hierarchy
In interaction modeling, practitioners often choose to enforce some form of hierarchy, which can be viewed as a type of regularization analogous to sparsity constraints. Two common forms of hierarchy are:
  • Strong hierarchy requires that if an interaction term $\widehat{\gamma}_{jk}\neq 0$, then both associated main effects $\widehat{\beta}_{j}$ and $\widehat{\beta}_{k}$ must also be nonzero.
  • Weak hierarchy only requires that at least one of the main effects is nonzero when an interaction is present i.e $\widehat{\gamma}_{jk}\neq 0\Rightarrow\widehat{\beta}_{j}\neq 0$ or $\widehat{\beta}_{k}\neq 0$.
The following two motivations are discussed in more details in Bien et al. [2]. First, a classical justification for enforcing strong hierarchy, as discussed by McCullagh and Nelder [8], is the following: suppose the model takes the form $Y=\beta_{0}+(\beta_{1}+\gamma_{12}X_{2})X_{1}+\ldots$ If $\beta_{1}=0$ while $\gamma_{12}\neq 0$ and under the assumption that there is no reason to distinguish $X_{2}$ over $X_{2}+c$ from some non-zero constant $c$, then the model with $X_{2}+c$ instead of $X_{2}$ is strongly hierarchical. Second, another justification, offered by Cox [4], is that large main effects are more likely to lead to significant interactions than small ones.
In Lim and Hastie [7], the authors distinguish between parameter sparsity and practical sparsity. The first being number of non-zero coefficients while the latter is the number of raw features used to make a prediction. Consider $Y=X_{1}+X_{2}+X_{1}X_{2}$ then parameter sparsity is $3$ while practical sparsity is $2$. So for a given value of parameter sparsity, a strongly hierarchical model has smaller practical sparsity than a non-hierarchical method.
4.2 Sprinter
In Yu et al. [12], the main guiding principle is that "one should prefer main effects over interactions if all else is equal". This assumption is weaker than hierarchy constraints but still introduces a bias in favor of including main effects unless there is strong evidence supporting interactions. In particular, an interaction is kept only if it can not be explained by a linear combination of main effects.
The paper introduces Sprinter which works as follows :
  1. 1.
    Fit a Lasso model on the main effects. Let $\hat{\theta}$ be a solution to
[…]
  2. 2.
    Compute residuals $R=Y-X\hat{\theta}$ and screen for interactions using residual correlation
    $\widehat{\Gamma}=\left\{(j,k)\in[p]^{2}:j<k,\;\widehat{\mathrm{SD}}(R)|\widehat{\mathrm{Corr}}(Z_{jk},R)|>\eta\right\}$
    where $\widehat{\mathrm{SD}}$(resp. $\widehat{\mathrm{Corr}}$) is the empirical standard deviation (resp. correlation). This step is a form of Sure Independence Screening (SIS) introduced in Fan and Lv [5].
  3. 3.
[…]
The final predictive model for any input $x\in\mathbb{R}^{p}$ is given by $x^{\top}(\widehat{\theta}+\widehat{\xi})+z_{\widehat{\Gamma}}^{\top}\widehat{\phi}$.
If an interaction is highly correlated with a main effect, then APL may select the interaction while Sprinter by design will prioritize the main effect instead.
For comparison, Interaction Pursuit (IP), as introduced in Fan et al. [6], first selects a set of strong main effects and then considers only the interactions among them. This enforces strong hierarchy by construction.
To avoid a three-dimensional cross-validation on $(\lambda_{1},\eta,\lambda_{3})$ in Sprinter, the authors propose the following : instead of thresholding residual-correlation by a fixed value $\eta$ define
$\widehat{\Gamma}=\left\{(j,k)\in[p]^{2}:j<k,\;|\widehat{\mathrm{Corr}}(Z_{jk},R)|\text{ is among the top }m\text{ largest}\right\}$
where $m\sim\frac{n}{\log n}$. Then a path algorithm is applied for two-dimensional cross-validation over $(\lambda_{1},\lambda_{3})$.
An extension, Sprinter+, performs cross-validation over $\lambda_{1}$ prior to step 2. This variant is more computationally efficient but may suffer in terms of out-of-sample performance, particularly when the main effects are weak or absent.
4.3 Glinternet
In Lim and Hastie [7], the authors define an interaction between variables $x$ and $y$ in a function $f$ as being present if $f(x,y)$ cannot be decomposed into the sum of two univariate functions, i.e., $f(x,y)\neq g(x)+h(y)$ for any functions $g$ and $h$. The paper introduces Glinternet which formulates the problem as a group lasso optimization involving $p+\binom{p}{2}$ groups : one group for each main effect and one for each pairwise interaction. For continuous variables, Glinternet solves :
[…]
Glinternet enforces strong hierarchy almost surely through a group lasso penalty structure. Generically, if an interaction term enters the model then the corresponding main effects are also active.
In practice, Glinternet is scalable and can handle problems with tens of thousands of variables as evidenced by the GWAS example in [7] with $26801$ variables and $3500$ observations.
4.4 HierNet
In Bien et al. [2], the authors introduce HierNet which in the gaussian case solves the following convex optimization problem
[…]
$\sum_{k\neq j}|\Theta_{jk}|$ $\leq\beta_{j}^{+}+\beta_{j}^{-},\quad\forall j\in[p],$
$\beta_{j}^{+},\,\beta_{j}^{-}$ $\geq 0,\quad\forall j\in[p]$
HierNet is a convex relaxation of the All Pairs Lasso (APL) with the strong hierarchy constraint $\|\Theta_{j}\|_{1}\leq|\beta_{j}|$, which is non-convex. It is equivalent to solving:
[…]
which shows that it uses the hierarchical penalty $\max(|\beta_{j}|,\|\Theta_{j}\|_{1})$ instead of the standard group-lasso penalty $\|(\beta_{j},\Theta_{j})\|_{2}$
If the symmetry constraint $\Theta=\Theta^{T}$ is removed, the resulting model enforces generically weak hierarchy. At a fixed level of parameter sparsity, HierNet typically achieves lower practical sparsity than APL and so uses fewer raw features for prediction. According to Lim and Hastie [7], HierNet works in practice for problems with up to $p<1000$ features.
4.5 Group Lasso Approaches
As discussed in Bien et al. [2], one way to enforce hierarchical structure in interaction modeling is to apply Lasso regularization over both main effects and interaction terms using a penalty like
$\sum_{1\leq j<k\leq p}|\Theta_{jk}|+\|(\beta_{j},\beta_{k},\Theta_{jk})\|_{2}$
This encourages sparsity at both the interaction coefficient level through $|\Theta_{jk}|$ and the group level through the $\ell_{2}$ norm across the group of interaction and corresponding main effects.
The idea is motivated by a general principle: a penalty of the form $\|(\beta_{i},\beta_{j})\|_{2}+\|\beta_{j}\|_{2}$ induces a hierarchical dependence of $\beta_{j}$ on $\beta_{i}$. Specifically, under this penalty, the condition $\beta_{j}\neq 0\Rightarrow\beta_{i}\neq 0$ is generically enforced. However, the converse does not necessarily hold i.e $\beta_{j}=0$ does not imply $\beta_{i}=0$.
5 More on the HIV mutation example
Recall the HIV mutation example of Section 2.3, where we saw that the uniPairs algorithms found interactions that display strong marginal interaction effects.
Figure 2: Marginal interactions for all feature pairs (grey), the pairs found by Glinternet (blue) and uniPairs-2stage (green). The interactions are measured by $\widehat{\beta}_{jk,jk}$ from the pairwise model (1).
Figure 2 compares the marginal interactions found by Glinternet and uniPairs with the marginal interactions for all pairs. We see that Glinternet pairs display fairly average marginal interactions, while the marginal interactions from uniPairs-2stage are in the positive tail of the distribution.
Figure 3 shows results from four methods applied to $50$ train/test draws of the HIV mutation data. We see that the uniPairs procedures have slightly higher (Test) MSE than Glinternet but win in every other measure.
Figure 3: Results from 50 train/test draws of the HIV mutation data. The “stability” is the average number features shared by each pair of 50 simulations. The Lasso baseline is fit using only main-effects, and therefore can’t select interaction terms.
6 A Simulation study
6.1 Simulation settings
Below we compare the performance of uniPairs-2stage, uniPairs, Sprinter(1cv), Glinternet on simulated data. HierNet was too slow to run. We adopt the data generating process described in section 5.2 of Yu et al. [12], omitting squared effects. More precisely, let $X\in\mathbb{R}^{n\times p}$ denote the design matrix of $p$ features. Each row of $X$ is independently sampled from a multivariate Gaussian distribution with zero mean and an AR(1) covariance structure $\mathrm{Cov}(X_{i},X_{j})=\rho^{|i-j|}$ where $\rho\in[0,1)$. The data-generating mechanism depends on a specified structure which determines the sets of active main effects $T_{1}\subseteq[p]$ and active interaction pairs $T_{3}\subseteq\{(j,k):1\leq j<k\leq p\}$. The considered structures are:
  • Mixed: both main effects and interactions are present without structural constraints.
    $T_{1}=\{0,1,2,3,4,5\}$ and $T_{3}=\{(0,4),(3,17),(9,10),(8,16),(0,12),(3,16)\}$
  • Hierarchical: interactions respect weak hierarchy.
    $T_{1}=\{0,1,2,3,4,5\}$ and $T_{3}=\{(0,2),(1,3),(2,3),(0,7),(1,7),(4,9)\}$
  • Anti-hierarchical: interactions occur only between features with no main effects (violating weak hierarchy).
    $T_{1}=\{0,1,2,3,4,5\}$ and $T_{3}=\{(10,12),(11,13),(12,13),(10,17),(11,17),(14,19)\}$
  • Interaction-only: only pairwise interactions are active.
    $T_{1}=\varnothing$ and $T_{3}=\{(0,2),(1,3),(2,3),(0,7),(1,7),(4,9)\}$
  • Main-effects-only: only main effects are active.
    $T_{1}=\{0,1,2,3,4,5\}$ and $T_{3}=\varnothing$
A coefficient vector $\beta\in\mathbb{R}^{p}$ is defined by $\beta_{j}=0$ unless $j\in T_{1}$ in which case $\beta_{j}=2$. The main effect signal is then $\mu_{\text{main}}=X\beta$. For the active interactions $(j,k)\in T_{3}$, the interaction signal is $\mu_{\text{interact}}=\sum_{(j,k)\in T_{3}}3\,(X_{\cdot j}\odot X_{\cdot k})$
The interaction component is orthogonalized with respect to the column space of the active main-effect features. Let $F=X_{[:,T_{1}]}$ and define the projection matrix $P_{F}=F(F^{\top}F)^{-1}F^{\top}$. Then, we do
$\mu_{\text{main}}\leftarrow\mu_{\text{main}}+P_{F}\mu_{\text{interact}}\quad\text{and}\quad\mu_{\text{interact}}\leftarrow\mu_{\text{interact}}-P_{F}\mu_{\text{interact}}$
After orthogonalization, the interaction component is rescaled to match the variance of the main component: $\mu_{\text{interact}}\leftarrow\mu_{\text{interact}}\times\sqrt{\frac{\mathrm{Var}(\mu_{\text{main}})}{\mathrm{Var}(\mu_{\text{interact}})}}$. The total signal is then $\mu=\mu_{\text{main}}+\mu_{\text{interact}}$.
Gaussian noise $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ is added, where the noise variance is set to achieve the desired signal-to-noise ratio $\sigma^{2}=\frac{\mathrm{Var}(\mu)}{\text{SNR}}$. The observed response is then $Y=\mu+\varepsilon$.
Our evaluation metrics are :
  • Test $R^{2}$ and Train $R^{2}$
  • Coverage : The fraction of true active variables correctly identified among the true actives, computed separately for main effects, interactions, and jointly.
  • False Discovery Rate (FDR) : The fraction of falsely selected variables among all predicted actives, computed for main effects, interactions, and both combined.
  • Model size : The total number of selected active variables, reported for main effects, interactions, and overall.
We consider four algorithms, five data-generating structures, and three SNRs $[0.5,1,3]$. The number of features and samples varies across configurations according to
$(n,p)\in\{(1000,80),(100,80),(300,400),(100,400),(100,200),(300,200)\}$
The correlation between features follows an AR(1) with $\rho\in\{0.0,0.25,0.5,0.75,1.0\}$. For each combination of structure, SNR, $\rho$, and $(n,p)$ pair, we perform 40 independent simulation replicates.
6.2 Simulation results
Method Glinternet Sprintr uniPairs uniPairs-2stage
Test $R^{2}$ 2.20 2.69 2.87 2.18
Train $R^{2}$ 2.64 2.14 2.76 2.39
Coverage Both 1.58 3.59 2.48 2.29
Coverage Main 1.92 3.34 2.46 2.23
Coverage Interactions 1.73 3.17 2.52 2.52
FDP Both 3.06 3.35 1.75 1.82
FDP Main 3.50 2.31 1.92 2.23
FDP Interactions 2.30 3.26 2.22 2.17
Model size Both 3.35 2.90 1.73 1.99
Model size Main 3.61 1.98 1.98 2.38
Model size Interactions 3.09 3.21 1.87 1.81
Table 1: Global average rank of each method across all simulations. Lower is better across all metrics.
In Table 1, we see that both uniPairs-2stage and Glinternet attain the best average rank in Test $R^{2}$. uniPairs comes next, then Sprinter. Across coverage metrics, Glinternet ranks best. However, this comes at a substantial cost: Glinternet ranks worst on FDR and model size, confirming that its high coverage is achieved through aggressive over-selection. In contrast, uniPairs and uniPairs-2stage achieve the top ranks for FDR, both overall and separately for main and interaction terms. The two variants deliver the smallest model sizes. They identify both main effects and interactions more conservatively leading to more interpretable models.
Arrows indicate direction of improvement: $\uparrow$ = larger is better,  $\downarrow$ = smaller is better.
Method Glinternet Sprintr uniPairs uniPairs-2stage
Test $R^{2}\uparrow$ 0.35 0.33 0.30 0.32
Train $R^{2}\uparrow$ 0.50 0.52 0.49 0.51
Coverage Both  $\uparrow$ 0.73 0.32 0.49 0.51
Coverage Main  $\uparrow$ 0.66 0.43 0.54 0.57
Coverage Interactions  $\uparrow$ 0.49 0.06 0.21 0.21
FDP Both  $\downarrow$ 0.79 0.83 0.51 0.54
FDP Main  $\downarrow$ 0.79 0.40 0.33 0.44
FDP Interactions  $\downarrow$ 0.71 0.91 0.60 0.59
Model size Both  $\downarrow$ 5924.84 27.94 10.27 11.45
Model size Main  $\downarrow$ 66.84 7.26 6.32 7.60
Model size Interactions  $\downarrow$ 5858.00 20.68 3.95 3.85
Table 2: The average value of each metric for each method across all simulations.
Table 2 confirms the patterns observed in Table 1. Both uniPairs and uniPairs-2stage provide the best balance between predictive performance and model parsimony. Coverage is highest for Glinternet across both main effects and interactions but this comes with the cost of substantially producing very large models with an average of nearly $6000$ selected terms compared to only 101210-12 for the uniPairs methods.
Arrows indicate direction of improvement: $\uparrow$ = larger is better,  $\downarrow$ = smaller is better.
Method Sprintr uniPairs uniPairs-2stage
Test $R^{2}\downarrow$ 0.02 0.06 0.03
Train $R^{2}\downarrow$ -0.04 0.16 0.12
Coverage Both  $\uparrow$ -0.82 -0.48 -0.43
Coverage Main  $\uparrow$ -0.44 -0.26 -0.20
Coverage Interactions  $\uparrow$ -1.32 -0.70 -0.70
FDP Both  $\downarrow$ 0.09 -0.43 -0.41
FDP Main  $\downarrow$ -0.42 -0.51 -0.43
FDP Interactions  $\downarrow$ 0.15 -0.11 -0.12
Model size Both  $\downarrow$ -1.48 -2.26 -2.10
Model size Main  $\downarrow$ -1.70 -1.55 -1.32
Model size Interactions  $\downarrow$ -1.26 -2.96 -2.93
Table 3: Log relative performance of each method compared to Glinternet across all simulations. For Test $R^{2}$ and Train $R^{2}$, the reported values are $\bm{\log\!\big[(1-R^{2}_{\texttt{method}})/(1-R^{2}_{\texttt{Glinternet}})\big]}$ so that a negative value indicate a reduction in squared error relative to Ginternet. For all other metrics $\bm{M}$, the reported values are $\bm{\log(M_{\texttt{method}}/M_{\texttt{Glinternet}})}$ with negative (resp. positive) values indicating an improvement over Glinternet in FDP and Model size (resp. Coverage). All cases where the logarithm is ill-defined are removed.
In Table 3, we see that both uniPairs and uniPairs-2stage achieve substantially lower model sizes without sacrificing predictive performance. Models from the uniPairs methods are roughly $5\%$ the size of the Glinternet model while Test $R^{2}$ values are slightly negative. Coverage decreases relative to Glinternet, as expected. However, the loss in coverage is modest and is offset by the improvements in sparsity and FDR. Sprinter shows weaker improvements. It reduces model size relative to Glinternet, but not as strongly as the uniPairs variants.
The next four figures show the overall model size, Test $R^{2}$, overall coverage and overall FDR for the case $n=300,p=400$. The other metrics are given in Appendix Appendix C: Full simulation results.
Figure 4: Total model size (main + interaction) for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$. The red dashed line marks the true number of active effects. Glinternet is plotted against the right y-axis while uniPairs, uniPairs-2stage and Sprinter use the left y-axis.
In Figure 4, we see that across all data-generating structures and SNR levels, uniPairs and uniPairs-2stage produce sparse models whose sizes are consistently close to the true number of active variables. The two variants behave similarly with uniPairs yielding slightly smaller models on average. Sprinter selects substantially more variables than needed and Glinternet selects orders of magnitude more terms than the true model, independent of structure and SNR level. Model size patterns remain stable across the different configurations.
Figure 5: Test $R^{2}$ for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 5, we see that Test $R^{2}$ increases with SNR as expected. uniPairs and uniPairs-2stage consistently achieve near-highest Test $R^{2}$ with uniPairs-2stage slightly performing better than uniPairs on average. Glinternet exhibits competitive Test $R^{2}$ overall, but not much larger than uniPairs and uniPairs-2stage despite fitting far larger models as seen in Figure 4.
Figure 6: Overall coverage (fraction of true active variables correctly identified) for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 6, we see that Glinternet achieves the highest coverage across all structures and SNR levels. This reflects Glinternet agressive selection behavior : it includes nearly all true active variables but at the cost of extremely large model sizes. uniPairs and uniPairs-2stage have moderate coverage and perform similarly across all structures. Sprinter exhibits the lowest coverage overall.
Figure 7: Overall false discovery rate (FDR) for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across 200+ replicates.
In Figure 7, we see that both Glinternet and Sprinter exhibit high overall FDR, even for high SNRs. uniPairs and uniPairs-2stage achieve lower overall FDR, with small changes across structures and SNR levels.
6.3 Summary of the simulation results
First, across a wide range of $n,p$, correlation levels $\rho$, SNRs and structures (mixed, hierarchical, anti-hierarchical, interaction-only, and main-effects-only), both uniPairs and uniPairs-2stage achive Test $R^{2}$ that is competitive with Glinternet and Sprinter as seen in Figure 5. This advantage is more pronounced at low SNRs where overfitting is a concern.
Second, uniPairs and uniPairs-2stage tend to produce substantially smaller predictive models in terms of the number of selected main effects and interactions as seen in Figures 4 8 9. In many settings, they are at the same level in Test $R^{2}$ as Glinternet while using noticeably fewer interactions terms (often orders of magnitude less), which leads to better practical sparsity.
Third, uniPairs and uniPairs-2stage maintain a very low FDR compared to Glinternet and Sprinter as seen in Figures 7 11 12. Their coverage exceeds that of Sprinter and is below that of Glinternet as seen in Figures 6 13 14. It is worth noting, however, that Glinternet attains its higher coverage while producing models that are typically far larger than those of the other three methods.
7 Theoretical results
7.1 The UniLasso step
In this section, we give theoretical guarantees for the main-effects UniLasso step used in uniPairs-2stage, and outline how the result can be extended to uniPairs. At a high level, the main result (Theorem 1) shows that under standard assumptions and for a suitable choice of the regularization penalty $\lambda_{1}$, with probability tending to one as $n$ goes to $\infty$, the UniLasso estimator in the first stage of uniPairs-2stage doesn’t select spurious main effects outside the true support and the estimated main effects coefficients are uniformly close to the true coefficients with error $O(\lambda_{1})$. Compared to the results in Chatterjee et al. [3], the rates are adapted to the presence of interaction terms and the noise level now contains both the contribution from omitted interactions and the marginal effects of non-active features. We also discuss how a similar argument applies to uniPairs (Theorem 2) when the TripletScan selected set is treated as a fixed set that contains all truly active interactions. We give the full proofs in Appendix Appendix A: Proofs.
Theorem 1
Consider i.i.d observations from the data-generating model
$Y\;=\;\beta^{*}_{0}1+X\beta^{\star}+Z\gamma^{\star}+\varepsilon$
with sparse supports $S_{M}=\mathrm{supp}(\beta^{\star})$ and $S_{I}=\mathrm{supp}(\gamma^{\star})$. The following result is about the coefficients of the UniLasso step in uniPairs-2stage without the initial standardization of main-effects. We follow the notation in 3.3 with no standardization, so $\widetilde{X}_{ij}$ is defined as $X_{ij}$. Assume
  1. (A1)
    $X_{i}$ is sub-gaussian with $\|X_{i}\|_{\Psi_{2}}<C_{1}<\infty$ where $C_{1}$ is a positive absolute constant.
  2. (A2)
    $\epsilon_{i}$ is sub-exponential with $\|\epsilon_{i}\|_{\Psi_{1}}<C_{2}<\infty$ where $C_{2}$ is a positive absolute constant, and $\mathbb{E}(\epsilon_{i})=0$.
  3. (A3)
    Let $S_{M}^{\prime}=S_{M}\cup\{0\}$ and define $X_{i0}=1$. Let
    $\Sigma_{S_{M}^{\prime}}=\mathbb{E}\big[X_{i,S_{M}^{\prime}}X_{i,S_{M}^{\prime}}^{\top}\big]\quad\text{and}\quad\eta^{*}_{M}=\lambda_{\min}(\Sigma_{S_{M}^{\prime}})$
    Assume $\eta^{*}_{M}>c_{3}$ and $\|\Sigma_{S_{M}^{\prime}}\|_{op}<C_{4}$ where $0<c_{3},C_{4}<\infty$ are absolute constants.
  4. (A4)
    For each $j\in[p]$ define
    $\beta_{1,j}^{*,\mathrm{uni}}=\frac{\mathrm{Cov}(X_{ij},Y_{i})}{\mathrm{Var}(X_{ij})}\quad\text{and}\quad\beta_{0,j}^{*,\mathrm{uni}}=\mathbb{E}[Y_{i}]-\beta_{1,j}^{*,\mathrm{uni}}\mathbb{E}[X_{ij}]$
    Assume that for all $j\in S_{M}$,
    $\beta_{1,j}^{*,\mathrm{uni}}\beta_{j}^{*}>0$
    and for all $j\in[p]$, $\mathrm{Var}(X_{ij})>c_{5}$ and $|\beta_{1,j}^{*,\mathrm{uni}}|>c_{6}$ where $0<c_{5},c_{6}<\infty$ are positive absolute constants.
  5. (A5)
    Let
    $\beta_{0}^{*}=\beta_{0,M}^{*}+\beta_{0,I}^{*}\quad\text{where}\quad\beta_{0,M}^{*}=\mathbb{E}[Y_{i}]-\sum_{k\in S_{M}}\beta_{k}^{*}\mathbb{E}[X_{ik}]$
    and define
    $\varepsilon_{i}^{\prime}=Y_{i}-\beta_{0,M}^{*}-\sum_{k\in S_{M}}\beta_{k}^{*}X_{ik}$
    Let
    $B=\max\big\{\max_{k\in S_{M}}\big|\mathbb{E}[\varepsilon_{i}^{\prime}X_{ik}]\big|,\max_{j\notin S_{M}}|\beta_{1,j}^{*,\mathrm{uni}}|\big\}$
    Assume that
    $c_{7}B\leq\lambda_{1}\leq C_{8}$
    for some absolute constants $0<c_{7},C_{8}<\infty$.
  6. (A6)
    Assume that as $n\to\infty$
    $\log(pn)=o\!\big(n^{3/5}\lambda_{1}^{2}\big)\quad\text{and}\quad n^{3/5}\lambda_{1}^{2}\to\infty$
  7. (A7)
    Assume that $|S_{M}|$, $|S_{I}|$, $|\beta_{0,M}^{*}|$, $|\beta_{0,I}^{*}|$, $\max_{j\in S_{M}^{\prime}}|\beta_{j}^{*}|$, $\max_{(j,k)\in S_{I}}|\gamma_{jk}^{*}|$, $\mathbb{E}(\epsilon_{i}^{2})$ and $\max_{j\in[p]}\mathbb{E}(X_{ij}^{4})$ are all upper bounded by a positive absolute constant $0<C_{9}<\infty$.
Then there exists absolute constants $C,c>0$ depending only on the absolute quantities in the assumptions such that for all $n$ large enough,
$\mathbb{P}\Big(\forall j\notin S_{M}^{\prime}\ \widehat{\beta}_{j}^{s}=0,\max_{j\in S_{M}}|\widehat{\beta}_{j}^{s}-\beta^{*}_{j}|\leq C\lambda_{1},|\widehat{\beta}_{0}^{s}-\beta^{*}_{0,M}|\leq C\lambda_{1}\Big)$
$\geq 1-Cpn\,\exp(-cn^{3/5}\lambda_{1}^{2})-Cn\,\exp(-cn^{3/5})\xrightarrow[n\to\infty]{}1$
The main differences with the result in Chatterjee et al. [3] are
  1. 1.
    The convergence rates $Cpn\,\exp(-cn^{3/5}\lambda_{1}^{2})+Cn\,\exp(-cn^{3/5})$ compared with $Cpn\,\exp(-cn\lambda_{1}^{2})$.
  2. 2.
    The noise level
    $\max\big\{\max_{k\in S_{M}}\big|\mathbb{E}[\varepsilon_{i}^{\prime}X_{ik}]\big|,\max_{j\notin S_{M}}|\beta_{1,j}^{*,\mathrm{uni}}|\big\}=O\!\big(\lambda_{1}\big)$
    compared with $\max_{j\notin S_{M}^{\prime}}|\beta_{1,j}^{*,\mathrm{uni}}|=O\big(\lambda_{1}\big)$.
Next, we provide an extension of Theorem 1 that deals with the coefficients of uniPairs assuming that the TripletScan selected set is a fixed set that contains all the true active interactions. Without the previous assumption, the TripletScan selected set depends on samples $(X_{i},Y_{i})$ and so is itself a random quantity.
Theorem 2
Consider i.i.d observations from the data-generating model
$Y\;=\;\beta^{*}_{0}1+X\beta^{\star}+Z\gamma^{\star}+\varepsilon$
with sparse supports $S_{M}=\mathrm{supp}(\beta^{\star})$ and $S_{I}=\mathrm{supp}(\gamma^{\star})$. The following result is about the coefficients of the UniLasso step in uniPairs without the initial standardization of main-effects and assuming that the TripletScan selected set is a fixed set that contains all the true active interactions. We follow the notation in 3.2 with no standardization, so $\widetilde{X}_{ij}$ is defined as $X_{ij}$. Define $X^{\mathrm{aug}}=[X|Z]\in\mathbb{R}^{n\times\left(p+\binom{p}{2}\right)}$.
Suppose Assumptions A1 A2 A4 of Theorem 1 hold. In addition, assume :
  1. (A8)
    Let $S_{A}=S_{M}\cup S_{I}$ and $S_{A}^{\prime}=S_{A}\cup\{0\}$ and define
    $\Sigma_{S_{A}^{\prime}}=\mathbb{E}\big[X_{i,S_{A}^{\prime}}^{\mathrm{aug}}X_{i,S_{A}^{\prime}}^{\mathrm{aug}\top}\big]\quad\text{and}\quad\eta^{*}_{A}=\lambda_{\min}(\Sigma_{S_{A}^{\prime}})$
    Assume $\eta^{*}_{A}>c_{10}$ and $\|\Sigma_{S_{A}^{\prime}}\|_{op}<C_{11}$ where $0<c_{10},C_{11}<\infty$ are absolute constants.
  2. (A9)
    For each $(j,k)\in\mathcal{P}$ define
    $\beta_{1,jk}^{*,\mathrm{uni}}=\frac{\mathrm{Cov}(X_{ij}X_{ik},Y_{i})}{\mathrm{Var}(X_{ij}X_{ik})}\quad\text{and}\quad\beta_{0,jk}^{*,\mathrm{uni}}=\mathbb{E}[Y_{i}]-\beta_{1,j}^{*,\mathrm{uni}}\mathbb{E}[X_{ij}X_{ik}]$
    Assume that for all $(j,k)\in S_{I}$,
    $\beta_{1,jk}^{*,\mathrm{uni}}\gamma_{jk}^{*}>0$
    and for all $(j,k)\in\mathcal{P}$, $\mathrm{Var}(X_{ij}X_{ik})>c_{12}$ and $|\beta_{1,jk}^{*,\mathrm{uni}}|>c_{13}$ where $0<c_{12},c_{13}<\infty$ are positive absolute constants.
  3. (A10)
    $S_{I}=\{(j,k)\in\mathcal{P}\,|\,\gamma^{*}_{jk}\neq 0\}\subset\widehat{\Gamma}$ where $\widehat{\Gamma}$ is the set of selected interactions after the TripletScan step that is considered fixed.
  4. (A11)
    Let
    $B^{\prime}=\max\big\{\max_{(j,k)\notin S_{I}}|\beta_{1,jk}^{*,\mathrm{uni}}|,\max_{j\notin S_{M}}|\beta_{1,j}^{*,\mathrm{uni}}|\big\}$
    Assume that
    $c_{14}B^{\prime}\leq\lambda\leq C_{15}$
    for some absolute constants $0<c_{14},C_{15}<\infty$.
  5. (A12)
    Assume that as $n\to\infty$
    $\log(p^{2}n)=o\!\big(n^{3/5}\lambda^{2}\big)\quad\text{and}\quad n^{3/5}\lambda^{2}\to\infty$
  6. (A13)
    Assume that $|S_{M}|$, $|S_{I}|$, $\max_{j\in S_{M}^{\prime}}|\beta_{j}^{*}|$, $\max_{(j,k)\in S_{I}}|\gamma_{jk}^{*}|$, $\mathbb{E}(\epsilon_{i}^{2})$ and $\max_{j\in[p]}\mathbb{E}(X_{ij}^{4})$ are all upper bounded by a positive absolute constant $0<C_{16}<\infty$.
Then there exists absolute constants $C^{\prime},c^{\prime}>0$ depending only on the absolute quantities in the assumptions such that for all $n$ large enough,
$\mathbb{P}\Big(\forall j\notin S_{M}^{\prime}\ \widehat{\beta}_{j}^{s}=0,\forall(j,k)\notin S_{I}\ \widehat{\beta}_{jk}^{s}=0,\max_{j\in S_{M}^{\prime}}|\widehat{\beta}_{j}^{s}-\beta^{*}_{j}|\leq C^{\prime}\lambda,\max_{(j,k)\in S_{I}}|\widehat{\beta}_{jk}^{s}-\gamma^{*}_{jk}|\leq C^{\prime}\lambda\Big)$
$\geq 1-C^{\prime}p^{2}n\,\exp(-c^{\prime}n^{3/5}\lambda_{1}^{2})-C^{\prime}n\,\exp(-c^{\prime}n^{1/3})\xrightarrow[n\to\infty]{}1$
7.2 Triplet Scans and largest log-gap rule
For each interaction candidate $(j,k)\in\mathcal{P}$, we consider the local model
$Y=\beta_{0,jk}+\beta_{j,jk}X_{j}+\beta_{k,jk}X_{k}+\beta_{jk,jk}X_{j}\odot X_{k}+\varepsilon$ (3)
This corresponds exactly to Conditional Sure Independence Screening (CSIS) as in Barut et al. [1] with conditioning set
$X_{\mathcal{C}}=\big[\,\mathbf{1},\,X_{j},\,X_{k}\,\big]\quad\text{and candidate variable}\quad X_{j}\odot X_{k}$
In CSIS, the goal is to screen variables based on the Conditional Linear Covariance
$\mathrm{Cov}_{L}(Y,X_{j}\mid X_{\mathcal{C}})=\mathbb{E}\left[\big(Y-L(Y\mid X_{\mathcal{C}})\big)\big(X_{j}-L(X_{j}\mid X_{\mathcal{C}})\big)\right]$
where the Conditional Linear Expectation $L(Z\mid X_{\mathcal{C}})$ of any random variable $Z$ given a conditioning set $X_{\mathcal{C}}$ is defined as the $L^{2}$-projection of $Z$ onto the span of $\{1,X_{\mathcal{C}}\}$, ie
$L(Z\mid X_{\mathcal{C}})=a^{*}+b^{*\,\top}X_{\mathcal{C}}\quad\text{where}\quad(a^{*},b^{*})\in\arg\min_{a,b}\mathbb{E}\big[(Z-a-b^{\top}X_{\mathcal{C}})^{2}\big]$
The Conditional Linear Covariance $\mathrm{Cov}_{L}(Y,X_{j}\mid X_{\mathcal{C}})$ quantifies the remaining linear dependence between $Y$ and $X_{j}$ after conditioning on $X_{\mathcal{C}}$. In our context,
$\beta_{jk,jk}=0\quad\Longleftrightarrow\quad\mathrm{Cov}_{L}(Y,X_{j}\odot X_{k}\mid\mathbf{1},X_{j},X_{k})=0$
which is a direct application of Theorem $1$ in Barut et al. [1]. Hence, $\beta_{jk,jk}$ captures the residual linear association between $Y$ and the interaction $X_{j}\odot X_{k}$ after adjusting for the two corresponding main-effects.
Let $\widehat{\beta}_{jk,jk}$ be the estimated coefficient for $\beta_{jk,jk}$ in 3. CSIS uses $|\widehat{\beta}_{jk,jk}|$ as the screening statistic for the interaction $X_{j}\odot X_{k}$. The screening rule is defined by thresholding the magnitude of the conditional marginal coefficient, ie
$\widehat{S}_{\gamma}=\{(j,k)\in\mathcal{P}:|\widehat{\beta}_{jk,jk}|\geq\gamma\}$
for some threshold $\gamma>0$. Let $S_{I}^{L}\subset\mathcal{P}$ denote the set of pairs $(j,k)$ satisfying $\mathrm{Cov}_{L}(Y,X_{j}\odot X_{k}\mid\mathbf{1},X_{j},X_{k})\neq 0$. Then, under the standard CSIS signal strength condition, namely that
$\min_{(j,k)\in S_{I}^{L}}\big|\mathrm{Cov}_{L}(Y,X_{j}\odot X_{k}\mid\mathbf{1},X_{j},X_{k})\big|\geq c\,n^{-\kappa}\quad\text{for some}\quad\kappa<\tfrac{1}{2}\quad\text{and}\quad c>0,$
together with regularity conditions stated in Barut et al. [1], the CSIS procedure satisfies the sure screening property, ie
$\mathbb{P}(S_{I}^{L}\subset\widehat{S}_{\gamma})\xrightarrow[n\to\infty]{}1$
for a threshold $\gamma\asymp n^{-\kappa}$.
The hierarchy restriction (strong or weak) in uniPairs-2stage simply makes the conditioning sets data-driven—replacing $X_{\mathcal{C}}$ by $\mathbf{1},X_{j},X_{k}]$ only when $j$ or $k$ has been identified as a main effect.
To select a threshold among the $p$-values $\{p_{jk}\}$, we use the “largest log-gap” rule. Let the ordered $p$-values over the eligible set $\mathcal{E}$ be $p_{(1)}\leq\cdots\leq p_{(M)}$ and define $\ell_{r}=\log p_{(r)}$. The rule selects
$\widehat{r}\in\arg\max_{1\leq r<M}\ (\ell_{r+1}-\ell_{r})\quad\text{then}\quad\widehat{\Gamma}=\big\{(j,k)\in\mathcal{E}:p_{jk}\leq p_{(\widehat{r})}\big\}$
This thresholding mechanism is motivated by the standard two-group mixture model for $p$-values:
[…]
where $G_{\text{alt}}$ is stochastically smaller than the uniform distribution. Under the null, the spacings of $\ell_{r}=-\log p_{(r)}$ are approximately exponential and nearly homogeneous. When signals are present, the smallest $p$-values form a tight cluster near zero, followed by a large jump as one transitions to the null regime. The maximal log-gap $(\ell_{r+1}-\ell_{r})$ therefore estimates the boundary between signal and noise, analogous to detecting an “elbow” in the empirical $\log p$ curve.
8 Discussion
We have introduced uniPairs and uniPairs-2stage, two univariate guided procedures for learning sparse interaction models in high dimensions. Both methods use a data-driven screening rule without additional hyperparameters, and they leverage the UniLasso framework to combine univariate fits into a multivariate predictor. Empirically, the proposed methods have competitive predictive performance relative to existing methods such as Sprinter and Glinternet, while selecting fewer interaction terms and thus producing models that are easier to interpret. Theoretically, the UniLasso main-effects stage has support consistency and $\ell_{\infty}$ control in both methods under suitable conditions, extending the guarantees of Chatterjee et al. [3]. Moreover, the TripletScan uses a conditional sure independence screening mechanism.
Directions for future work are :
  • Extend uniPairs and uniPairs-2stage to more general feature engineering pipelines.
  • The TripletScan screening step is embarrassingly parallel, so exploring GPU and more efficient parallel/vectorized implementations could further improve scalability when $p$ is large.
  • Run the TripletScan step on a validation set in both uniPairs and uniPairs-2stage. We believe this change will make the two variants have better out-of-sample performance.
Both packages uniPairs and uniPairs-2stage are available on the PyPI repository https://pypi.org/project/uniPairs/. Install via pip install uniPairs. The full documentation is available at https://aymenecharghaoui.github.io/uniPairs/ while the Github repository is at https://github.com/AymenEcharghaoui/uniPairs.
Acknowledgements. We thank Trevor Hastie for helpful comments. R.T. was supported by the NIH (5R01EB001988-16) and the NSF (19DMS1208164).
Appendix A: Proofs
Below we give a full proof of the Theorems 1 and 2. We follow closely Chatterjee et al. [3]. Assume that the population data-generating model is
$Y\;=\;\beta^{*}_{0}1+X\beta^{\star}+Z\gamma^{\star}+\varepsilon$
with sparse supports $S_{M}=\mathrm{supp}(\beta^{\star})$ and $S_{I}=\mathrm{supp}(\gamma^{\star})$. Recall that
$\mathcal{P}=\{(j,k)\in[p]^{2}\ |\ \ j<k\}\quad\text{and}\quad Z=\big(X_{j}\odot X_{k}\big)_{(j,k)\in\mathcal{P}}\in\mathbb{R}^{n\times\binom{p}{2}}$
For any function $f:\{1,\ldots,n\}\to\mathbb{R}$ and $i\in[n]$, define
$P_{n}[f]=\frac{1}{n}\sum_{l=1}^{n}f(l)$
$L_{n,i}[f]=\frac{1}{n-1}\sum_{\begin{subarray}{c}l=1\\ l\neq i\end{subarray}}^{n}f(l)$
Then, for each $j\in[p]$, the leave-one-out univariate regression coefficients from the main-effect $X_{j}$ are
$\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}=\frac{L_{n,i}[X_{\ell j}Y_{\ell}]-L_{n,i}[Y_{\ell}]\,L_{n,i}[X_{\ell j}]}{L_{n,i}[X_{\ell j}^{2}]-\big(L_{n,i}[X_{\ell j}]\big)^{2}}\quad\text{and}\quad\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}=L_{n,i}[Y_{\ell}]-\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}\,L_{n,i}[X_{\ell j}]$
and the corresponding $i^{\text{th}}$ leave-one-out prediction is
$\widehat{\eta}_{j}^{(-i)}=\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}+\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}X_{ij}$
Also, for each $(j,k)\in\mathcal{P}$, the leave-one-out univariate regression coefficients from the interaction $X_{j}\odot X_{k}$ are
$\widehat{\beta}_{1,jk}^{(-i)\mathrm{uni}}=\frac{L_{n,i}[X_{\ell j}X_{\ell k}Y_{\ell}]-L_{n,i}[Y_{\ell}]\,L_{n,i}[X_{\ell j}X_{\ell k}]}{L_{n,i}[X_{\ell j}^{2}X_{\ell k}^{2}]-\big(L_{n,i}[X_{\ell j}X_{\ell k}]\big)^{2}}\quad\text{and}\quad\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}=L_{n,i}[Y_{\ell}]-\widehat{\beta}_{1,jk}^{(-i)\mathrm{uni}}\,L_{n,i}[X_{\ell j}X_{\ell k}]$
and the corresponding $i^{\text{th}}$ leave-one-out prediction is
$\widehat{\eta}_{jk}^{(-i)}=\widehat{\beta}_{0,jk}^{(-i)\mathrm{uni}}+\widehat{\beta}_{1,jk}^{(-i)\mathrm{uni}}X_{ij}X_{ik}$
In the following, we write $O_{a}(1)$ (resp. $\Omega_{a}(1)$) to denote any positive expression that is bounded above (resp. below) by a positive absolute constant.
Our first lemma gives concentration bounds of empirical moments.
Lemma 1
Assume A1, A2 and A7. Then, there exists a positive absolute constant $0<K<\infty$ such that for any indices $j,k,r,m\in[p]$, any $t>0$, and all $n\geq 1$, we have :
$\mathbb{P}\!\left(|P_{n}[X_{ij}]-\mathbb{E}[X_{ij}]|>t\right)\leq 2\exp\!\left(-K\,nt^{2}\right)$
$\mathbb{P}\!\left(|P_{n}[X_{ij}X_{ik}]-\mathbb{E}[X_{ij}X_{ik}]|>t\right)\leq 2\exp\!\left(-K\,n\,\min\{t,t^{2}\}\right)$
$\mathbb{P}\!\left(|P_{n}[X_{ij}X_{ik}X_{ir}]-\mathbb{E}[X_{ij}X_{ik}X_{ir}]|>t\right)\leq 2\exp\!\left(-K\,t\,n^{3/4}\right)$
$\mathbb{P}\!\left(|P_{n}[X_{ij}X_{ik}X_{ir}X_{im}]-\mathbb{E}[X_{ij}X_{ik}X_{ir}X_{im}]|>t\right)\leq 2\exp\!\left(-K\,n^{3/5}t^{4/5}\right)$
$\mathbb{P}\!\left(|P_{n}[Y_{i}]-\mathbb{E}[Y_{i}]|>t\right)\leq 2\exp\!\left(-K\,n\,\min\{t,t^{2}\}\right)$
$\mathbb{P}\!\left(|P_{n}[Y_{i}^{2}]-\mathbb{E}[Y_{i}^{2}]|>t\right)\leq 2\exp\!\left(-K\,n^{3/5}t^{4/5}\right)$
$\mathbb{P}\!\left(|P_{n}[X_{ij}Y_{i}]-\mathbb{E}[X_{ij}Y_{i}]|>t\right)\leq 2\exp\!\left(-K\,t\,n^{3/4}\right)$
$\mathbb{P}\!\left(|P_{n}[X_{ij}X_{ik}Y_{i}]-\mathbb{E}[X_{ij}X_{ik}Y_{i}]|>t\right)\leq 2\exp\!\left(-K\,n^{3/5}t^{4/5}\right)$
$\mathbb{P}\!\left(|P_{n}[\epsilon_{i}^{\prime}X_{ik}]-\mathbb{E}[\epsilon_{i}^{\prime}X_{ik}]|>t\right)\leq 2\exp\!\left(-K\,n^{3/4}t\right)$
$\mathbb{P}\!\left(|P_{n}[|\epsilon_{i}^{\prime}X_{ik}|]-\mathbb{E}[|\epsilon_{i}^{\prime}X_{ik}|]|>t\right)\leq 2\exp\!\left(-K\,n^{3/4}t\right)$
All bounds remain valid if $P_{n}$ is replaced by $L_{n,i}$ for any fixed $i\in[n]$.
Proof.
Under A2, we have $\mathbb{E}[\varepsilon_{i}]=0$, so
$Y_{i}-\mathbb{E}[Y_{i}]=X_{i}^{\top}\beta^{*}-\mathbb{E}[X_{i}]^{\top}\beta^{*}+\tfrac{1}{2}X_{i}^{\top}\Gamma^{*}X_{i}-\tfrac{1}{2}\mathbb{E}[X_{i}^{\top}\Gamma^{*}X_{i}]+\varepsilon_{i}$
where $\Gamma^{*}=(\gamma_{jk}^{*})_{j,k\in[p]}$ is a symmetric matrix of interaction coefficients. By the triangle inequality for the $\psi_{1}$ norm,
$\|Y_{i}-\mathbb{E}[Y_{i}]\|_{\psi_{1}}\leq\|(X_{i}-\mathbb{E}[X_{i}])^{\top}\beta^{*}\|_{\psi_{1}}+\Big\|X_{i}^{\top}\Gamma^{*}X_{i}-\mathbb{E}[X_{i}^{\top}\Gamma^{*}X_{i}]\Big\|_{\psi_{1}}+\|\varepsilon_{i}\|_{\psi_{1}}$
Under A1, we have $\|X_{i}\|_{\psi_{2}}<\infty.$ By Lemma 2.7.7, Exercise 2.7.10 and Example 2.5.8 in Vershynin [11], we get
$\|(X_{i}-\mathbb{E}[X_{i}])^{\top}\beta^{*}\|_{\psi_{1}}\leq O_{a}(1)\,\|X_{i}\|_{\psi_{2}}\,\|\beta^{*}\|_{2}$
Since $X_{i}$ is sub-Gaussian, each product $X_{ij}X_{ik}$ is sub-exponential with
$\|X_{ij}X_{ik}\|_{\psi_{1}}\leq\|X_{ij}\|_{\psi_{2}}\|X_{ik}\|_{\psi_{2}}\leq\|X_{i}\|_{\psi_{2}}^{2}$
Hence,
$\Big\|X_{i}^{\top}\Gamma^{*}X_{i}-\mathbb{E}[X_{i}^{\top}\Gamma^{*}X_{i}]\Big\|_{\psi_{1}}$ $=\Big\|2\sum_{(j,k)\in S_{I}}\gamma_{jk}^{*}\big(X_{ij}X_{ik}-\mathbb{E}[X_{ij}X_{ik}]\big)\Big\|_{\psi_{1}}$
$\leq 2\sum_{(j,k)\in S_{I}}|\gamma_{jk}^{*}|\,\big\|X_{ij}X_{ik}-\mathbb{E}[X_{ij}X_{ik}]\big\|_{\psi_{1}}$
$\leq O_{a}(1)\,\|X_{i}\|_{\psi_{2}}^{2}\,\|\Gamma^{*}\|_{L^{1}}$
where $\|\Gamma^{*}\|_{L^{1}}=\sum_{j<k}|\gamma_{jk}^{*}|$. Under A2, we have $\|\varepsilon_{i}\|_{\psi_{1}}<\infty.$ Therefore,
$\|Y_{i}-\mathbb{E}[Y_{i}]\|_{\psi_{1}}\leq O_{a}(1)\Big(\|X_{i}\|_{\psi_{2}}\|\beta^{*}\|_{2}+\|X_{i}\|_{\psi_{2}}^{2}\|\Gamma^{*}\|_{L^{1}}+\|\varepsilon_{i}\|_{\psi_{1}}\Big)<\infty$
Hence, $Y_{i}-\mathbb{E}[Y_{i}]$ is sub-exponential, and therefore $Y_{i}$ is sub-exponential as well. By Young’s inequality,
$|X_{ij}Y_{i}|^{2/3}\leq\frac{1}{3}X_{ij}^{2}+\frac{2}{3}|Y_{i}|$
Hence, for any $\xi>0$, by Jensen’s inequality,
$\mathbb{E}\!\left[\exp\!\left(\frac{|X_{ij}Y_{i}|^{2/3}}{\xi^{2/3}}\right)\right]\leq\frac{1}{3}\mathbb{E}\!\left[\exp\!\left(\frac{X_{ij}^{2}}{\xi^{2/3}}\right)\right]+\frac{2}{3}\mathbb{E}\!\left[\exp\!\left(\frac{|Y_{i}|}{\xi^{2/3}}\right)\right]$
Since $\|X_{ij}\|_{\psi_{2}}<\infty$ and $\|Y_{i}\|_{\psi_{1}}<\infty$, both expectations are finite for suitable $\xi$. Therefore,
$\|X_{ij}Y_{i}\|_{\psi_{2/3}}=\inf\Bigl\{\xi>0:\mathbb{E}\!\left[\exp\!\left(\frac{|X_{ij}Y_{i}|^{2/3}}{\xi^{2/3}}\right)\right]\leq 2\Bigr\}<\infty$
Hence, $X_{ij}Y_{i}$ is sub-Weibull(2/3). By replacing $Y_{i}$ with $\epsilon_{i}$ in the previous equation and since $\epsilon_{i}$ is sub-exponential, we get that $\epsilon_{i}X_{il}$ is also sub-Weibull(2/3). Similarly, because
$|X_{ij}X_{ik}X_{ir}|^{2/3}\leq\frac{1}{3}X_{ij}^{2}+\frac{1}{3}X_{ik}^{2}+\frac{1}{3}X_{ir}^{2},\qquad|X_{ij}X_{ik}X_{ir}X_{im}|^{1/2}\leq\frac{1}{4}X_{ij}^{2}+\frac{1}{4}X_{ik}^{2}+\frac{1}{4}X_{ir}^{2}+\frac{1}{4}X_{im}^{2}$
and
$|X_{ij}X_{ik}Y_{i}|^{1/2}\leq\frac{1}{4}X_{ij}^{2}+\frac{1}{4}X_{ik}^{2}+\frac{1}{2}|Y_{i}|$
we get that $X_{ij}X_{ik}X_{ir}$ is sub-Weibull(2/3), $X_{ij}X_{ik}X_{ir}X_{im}$ and $X_{ij}X_{ik}Y_{i}$ are sub-Weibull(1/2). Also, $Y_{i}^{2}$ is sub-Weibull(1/2) because $\|Y_{i}\|_{\psi_{1}}<\infty$, and $X_{ij}^{2}$ is sub-exponential because $\|X_{ij}\|_{\psi_{2}}<\infty$. $X_{i}$ is sub-gaussian so in particular is sub-Weibull(2/3) $\big(|X_{i}|^{2/3}\leq 1+|X_{i}|^{2}\big)$. Hence,
$\|\epsilon_{i}^{\prime}X_{il}\|_{\psi_{2/3}}$ $\leq\sum_{(j,k)\in S_{I}}|\gamma^{*}_{j,k}|\|X_{ij}X_{ik}X_{il}\|_{\psi_{2/3}}+|\beta^{*}_{0,I}||\|X_{il}\|_{\psi_{2/3}}+\|\epsilon_{i}X_{il}\|_{\psi_{2/3}}<\infty$
We can then apply known concentration inequalities for sub-gaussian, sub-Weibull and sub-exponential variables (see, e.g., Theorem 9 Yu et al. [12], Theorem 2.8.1 Vershynin [11] and Theorem 2.6.3 Vershynin [11]). For all $t>0$, the followings hold:
  • $X_{ij}X_{ik}X_{ir}$, $X_{ij}Y_{i}$, $\epsilon_{i}^{\prime}X_{il}$, $|\epsilon_{i}^{\prime}X_{il}|$ are sub-Weibull(2/3), so by Theorem 9 Yu et al. [12],
    $\mathbb{P}\!\left(\big|P_{m}[X_{ij}X_{ik}X_{ir}]-\mathbb{E}[X_{ij}X_{ik}X_{ir}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}/\|X_{ij}X_{ik}X_{ir}\|_{\psi_{2/3}}\right)$
    $\mathbb{P}\!\left(\big|P_{m}[X_{ij}Y_{i}]-\mathbb{E}[X_{ij}Y_{i}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}/\|X_{ij}Y_{i}\|_{\psi_{2/3}}\right)$
    $\mathbb{P}\!\left(\big|P_{n}[\epsilon_{i}^{\prime}X_{il}]-\mathbb{E}[\epsilon_{i}^{\prime}X_{il}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}/\|\epsilon_{i}^{\prime}X_{il}\|_{\psi_{2/3}}\right)$
    $\mathbb{P}\!\left(\big|P_{n}[|\epsilon_{i}^{\prime}X_{il}|]-\mathbb{E}[|\epsilon_{i}^{\prime}X_{il}|]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}/\|\epsilon_{i}^{\prime}X_{il}\|_{\psi_{2/3}}\right)$
  • $Y_{i}$ and $X_{ij}X_{ik}$ are sub-exponential, so by Bernstein’s inequality (Theorem 2.8.2 in Vershynin [11]),
    $\mathbb{P}\!\left(\big|P_{n}[Y_{i}]-\mathbb{E}[Y_{i}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,n\,\min\!\Big\{\tfrac{t^{2}}{\|Y_{i}\|_{\psi_{1}}^{2}},\tfrac{t}{\|Y_{i}\|_{\psi_{1}}}\Big\}\right)$
    $\mathbb{P}\!\left(\big|P_{n}[X_{ij}X_{ik}]-\mathbb{E}[X_{ij}X_{ik}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,n\,\min\!\Big\{\tfrac{t^{2}}{\|X_{ij}X_{ik}\|_{\psi_{1}}^{2}},\tfrac{t}{\|X_{ij}X_{ik}\|_{\psi_{1}}}\Big\}\right)$
  • $X_{ij}$ is sub-Gaussian, so by Hoeffding’s inequality (Theorem 2.6.3 in Vershynin [11]),
    $\mathbb{P}\!\left(\big|P_{n}[X_{ij}]-\mathbb{E}[X_{ij}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,nt^{2}/\|X_{ij}\|_{\psi_{2}}^{2}\right)$
  • $X_{ij}X_{ik}X_{ir}X_{im}$, $X_{ij}X_{ik}Y_{i}$ and $Y_{i}^{2}$ are sub-Weibull(1/2), so by Theorem 9 Yu et al. [12],
    $\mathbb{P}\!\left(\big|P_{n}[X_{ij}X_{ik}X_{ir}X_{im}]-\mathbb{E}[X_{ij}X_{ik}X_{ir}X_{im}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,n^{3/5}\,t^{4/5}/\|X_{ij}X_{ik}X_{ir}X_{im}\|_{\psi_{1/2}}^{4/5}\right)$
    $\mathbb{P}\!\left(\big|P_{n}[X_{ij}X_{ik}Y_{i}]-\mathbb{E}[X_{ij}X_{ik}Y_{i}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,n^{3/5}\,t^{4/5}/\|X_{ij}X_{ik}Y_{i}\|_{\psi_{1/2}}^{4/5}\right)$
    $\mathbb{P}\!\left(\big|P_{n}[Y_{i}^{2}]-\mathbb{E}[Y_{i}^{2}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,n^{3/5}\,t^{4/5}/\|Y_{i}^{2}\|_{\psi_{1/2}}^{4/5}\right)$
From A1, A2, and A7 and using our previous results that control the $\|.\|_{\psi}$ of the different quantities of interest, we see in particular that $\|\Gamma^{*}\|_{L^{1}},\|\beta^{*}\|_{2}$ are $O_{a}(1)$ and so are $\|X_{ij}X_{ik}X_{ir}\|_{\psi_{2/3}}$, $\|X_{ij}Y_{i}\|_{\psi_{2/3}}$, $\|\epsilon_{i}^{\prime}X_{il}\|_{\psi_{2/3}}$, $\|Y_{i}\|_{\psi_{1}}$, $\|X_{ij}X_{ik}\|_{\psi_{1}}$, $\|X_{ij}\|_{\psi_{2}}$, $\|X_{ij}X_{ik}X_{ir}X_{im}\|_{\psi_{1/2}}$, $\|X_{ij}X_{ik}Y_{i}\|_{\psi_{1/2}}$, and $\|Y_{i}^{2}\|_{\psi_{1/2}}$. Therefore, the concentration inequalities stated in Lemma 1 follow. $\square$
The next two lemmas are about the restricted eigenvalue concentrations. Denote by $\lambda_{min}(A)$ (resp. $\lambda_{max}(A)$) the smallest (resp. largest) eigenvalue of the symmetric matrix $A$.
Lemma 2
Assume A1, A3 and A7. Let
$\widehat{\eta}_{M}=\lambda_{min}\Big(P_{n}[X_{i,S_{M}^{\prime}}X_{i,S_{M}^{\prime}}^{T}]\Big)$
Then there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for any $\epsilon\in(0,1)$, and all $n$ sufficiently large, we have :
$\mathbb{P}\left(|\widehat{\eta}_{M}-\eta^{*}_{M}|>K_{1}n^{-\epsilon/2}\right)\leq K_{1}\exp\left(-K_{2}n^{1-\epsilon}\right)$
Proof.
For any vector $V\in\mathbb{R}^{p}$, define $V_{S_{M}^{\prime}}=(V_{i})_{i\in S_{M}^{\prime}}\in\mathbb{R}^{|S_{M}^{\prime}|}$. By Theorem 4.6.1 and Exercise 4.7.3 in Vershynin [11], we get that for all $u\geq 0$ with probability at least $1-2\exp(-u)$, we have :
$\|P_{n}[X_{i,S_{M}^{\prime}}X_{i,S_{M}^{\prime}}^{T}]-\Sigma_{S_{M}^{\prime}}\|_{op}$ $\leq O_{a}(1)\max\big(1,\frac{\|X_{i,S_{M}^{\prime}}\|_{\Psi_{2}}^{2}}{\eta_{M}^{*2}}\big)\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\frac{|S_{M}^{\prime}|+u}{n}\right)\|\Sigma_{S_{M}^{\prime}}\|_{op}$
Recall that for any two symmetric matrices $A$ and $B$, we have
$\lambda_{min}(A)\leq\lambda_{min}(B)+\lambda_{max}(A-B)\leq\lambda_{min}(B)+\|A-B\|_{op}$
Therefore, since $\|X_{i,S_{M}^{\prime}}\|_{\Psi_{2}}\leq\|X_{i}\|_{\Psi_{2}}=O_{a}(1)$ and by A1 and A3, we get that for all $u\geq 0$ with probability at least $1-2\exp(-u)$
$|\widehat{\eta}_{M}-\eta^{*}_{M}|\leq O_{a}(1)\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\frac{|S_{M}^{\prime}|+u}{n}\right)$
So for all $u>0$ such that $\frac{|S_{M}^{\prime}|+u}{n}<1$ with probability at least $1-2\exp(-u)$
$|\widehat{\eta}_{M}-\eta^{*}_{M}|\leq O_{a}(1)\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}$
By taking $u=n^{1-\epsilon}$ for $\epsilon\in(0,1)$ and noting that $|S_{M}^{\prime}|=O_{a}(1)$ from A7, we get the concentration bound stated in Lemma 2. $\square$
Lemma 3
Assume A1, A8 and A13. Let
$\widehat{\eta}_{A}=\lambda_{min}\Big(P_{n}[X_{i,S_{A}^{\prime}}^{\mathrm{aug}}X_{i,S_{A}^{\prime}}^{\mathrm{aug}\top}]\Big)$
Then there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for any $\epsilon\in(0,\frac{2}{3})$, and all $n\geq 1$, we have :
$\mathbb{P}\left(|\widehat{\eta}_{A}-\eta^{*}_{A}|>K_{1}n^{-\epsilon/2}\right)\leq K_{1}\exp\left(-K_{2}n^{(3-2\epsilon)/5}\right)$
Proof.
We can’t directly use Theorem 4.6.1 in Vershynin [11] as $X_{i,S_{A}^{\prime}}^{\mathrm{aug}}$ is no longer a sub-gaussian random vector. However, the proof of Theorem 4.6.1 in Vershynin [11] can be adapted to our case by using sub-Weibull concentration inequalities in Step 2. In the following we use a different argument. We have
$|\widehat{\eta}_{A}-\eta_{A}^{*}|\leq\|P_{n}[X_{i,S_{A}^{\prime}}^{\mathrm{aug}}X_{i,S_{A}^{\prime}}^{\mathrm{aug}\top}]-\Sigma_{S_{A}^{\prime}}\|_{op}\leq|S_{A}^{\prime}|\|P_{n}[X_{i,S_{A}^{\prime}}^{\mathrm{aug}}X_{i,S_{A}^{\prime}}^{\mathrm{aug}\top}]-\Sigma_{S_{A}^{\prime}}\|_{\max}$
where the last inequality follows because for any $m\times m$ symmetric matrix $M$, we have $\|M\|_{op}\leq m\|M\|_{\max}$ where $\|M\|_{\max}=\max_{1\leq i,j\leq m}|M_{ij}|$. Apart from the top-left value which is zero, the values of $P_{n}[X_{i,S_{A}^{\prime}}^{\mathrm{aug}}X_{i,S_{A}^{\prime}}^{\mathrm{aug}\top}]-\Sigma_{S_{A}^{\prime}}$ are of the form :
  • $P_{n}[X_{ij}]-\mathbb{E}(X_{ij})$ for some $j\in S_{M}$
  • $P_{n}[X_{ij}X_{ik}]-\mathbb{E}(X_{ij}X_{ik})$ for some $(j,k)\in S_{M}^{2}\cup S_{I}$
  • $P_{n}[X_{ij}X_{ik}X_{ir}]-\mathbb{E}(X_{ij}X_{ik}X_{ir})$ for some $(j,k,r)$ such that some permutation of it is in $S_{M}\times S_{I}$
  • $P_{n}[X_{ij}X_{ik}X_{ir}X_{im}]-\mathbb{E}(X_{ij}X_{ik}X_{ir}X_{im})$ for some $(j,k,r,m)$ such that some permutation of it is in $S_{I}\times S_{I}$
Therefore from the corresponding concentration inequalities in 1, we get that for all $t>0$
$\mathbb{P}(|\widehat{\eta}_{A}-\eta_{A}^{*}|>|S_{A}^{\prime}|t)$ $\leq 2|S_{M}|\exp(-\Omega_{a}(1)nt^{2})+2(|S_{M}|^{2}+|S_{I}|)\exp(-\Omega_{a}(1)n\min(t,t^{2}))$
$+12|S_{M}||S_{I}|\exp(-\Omega_{a}(1)n^{\frac{3}{4}}t)+48|S_{I}|^{2}\exp(-\Omega_{a}(1)n^{\frac{3}{5}}t^{\frac{4}{5}})$
By taking $t=n^{-\epsilon/2}$ for $\epsilon\in(0,\frac{2}{3})$ and noting that $|S_{M}|,|S_{I}|,|S_{A}^{\prime}|=O_{a}(1)$ from A13, we get the concentration bound stated in Lemma 3. $\square$
The next lemma is about the concentration of the univariate coefficients used in uniPairs-2stage.
Lemma 4
Assume A1,  A2,  A4 and  A7. Then, there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for all $j\in[p]$, $n\geq 1$, and $t\in(0,K_{1})$, we have :
$\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/4}t^{2})$
and
$\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/4}t^{2})$
Also, for all $j\in[p]$, $i\geq 1$, $n\geq i$ and $t\in(0,K_{1})$, we have :
$\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/4}t^{2})$
and
$\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/4}t^{2})$
Proof.
Fix $i\geq 1$ and $j\in[p]$. Define
$A_{n}=L_{n,i}[X_{\ell j}Y_{\ell}]-L_{n,i}[Y_{\ell}]\,L_{n,i}[X_{\ell j}]\quad\text{and}\quad B_{n}=L_{n,i}[X_{\ell j}^{2}]-\big(L_{n,i}[X_{\ell j}]\big)^{2}$
Also let
$a=\mathbb{E}[X_{ij}Y_{i}]-\mathbb{E}[X_{ij}]\,\mathbb{E}[Y_{i}]=\mathrm{Cov}(X_{ij},Y_{i})\quad\text{and}\quad b=\mathbb{E}[X_{ij}^{2}]-(\mathbb{E}[X_{ij}])^{2}=\mathrm{Var}(X_{ij})>0$
Then
$\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}=\frac{A_{n}}{B_{n}}\quad\text{and}\quad\beta_{1,j}^{*,\mathrm{uni}}=\frac{a}{b}$
We have
$|B_{n}-b|$ $\leq|L_{n,i}[X_{\ell j}^{2}]-\mathbb{E}[X_{ij}^{2}]|+\big|(L_{n,i}[X_{\ell j}])^{2}-(\mathbb{E}[X_{ij}])^{2}\big|$
$\leq|L_{n,i}[X_{\ell j}^{2}]-\mathbb{E}[X_{ij}^{2}]|+|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|(|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|+2\mathbb{E}[X_{ij}])$
$\leq|L_{n,i}[X_{\ell j}^{2}]-\mathbb{E}[X_{ij}^{2}]|+|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|^{2}+O_{a}(1)|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|$
Hence, using the concentration of $L_{n,i}[X_{\ell j}^{2}]$ and $L_{n,i}[X_{\ell j}]$ from Lemma 1, and since $b=\Omega_{a}(1)$ from A4, we get
$\mathbb{P}(|B_{n}-b|>b/2)\leq O_{a}(1)\exp(-\Omega_{a}(1)\,n)$
On $|B_{n}-b|\leq b/2$, we have $B_{n}\geq b/2=\Omega_{a}(1)$. Decompose
$A_{n}-a=\left(L_{n,i}[X_{\ell j}Y_{\ell}]-\mathbb{E}[X_{ij}Y_{i}]\right)-\left(L_{n,i}[Y_{\ell}]L_{n,i}[X_{\ell j}]-\mathbb{E}[Y_{i}]\mathbb{E}[X_{ij}]\right)$
But,
$|L_{n,i}[Y_{\ell}]L_{n,i}[X_{\ell j}]-\mathbb{E}[Y_{i}]\mathbb{E}[X_{ij}]|$ $\leq|L_{n,i}[Y_{\ell}]|\,|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|+|\mathbb{E}[X_{ij}]|\,|L_{n,i}[Y_{\ell}]-\mathbb{E}[Y_{i}]|$
So by A7,
$|L_{n,i}[Y_{\ell}]L_{n,i}[X_{\ell j}]-\mathbb{E}[Y_{i}]\mathbb{E}[X_{ij}]|$ $\leq O_{a}(1)\Big(|L_{n,i}[Y_{\ell}]-\mathbb{E}[Y_{i}]|+|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|$
$+$ $|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]||L_{n,i}[Y_{\ell}]-\mathbb{E}[Y_{i}]|\Big)$
Using the concentration inequalities for each term from Lemma 1, we see that for $t\in(0,1)$
$\mathbb{P}\big(|A_{n}-a|>t\big)\leq O_{a}(1)\,\exp(-\Omega_{a}(1)\,n^{3/4}t^{2})$
But,
$\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{\star}=\frac{A_{n}-a}{B_{n}}+\frac{a}{b}\frac{b-B_{n}}{B_{n}}$
Therefore on $|B_{n}-b|\leq b/2$, we have
$|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|\leq O_{a}(1)\,|A_{n}-a|+O_{a}(1)\,|B_{n}-b|$
Therefore, we get
$\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|>t\right)\leq O_{a}(1)\,\exp(-\Omega_{a}(1)\,n^{3/4}t^{2})+O_{a}(1)\,\exp(-\Omega_{a}(1)\,n)$
Since $t\in(0,1)$, the second term is dominated by the first, yielding
$\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|>t\right)\leq O_{a}(1)\,\exp(-\Omega_{a}(1)\,n^{3/4}t^{2})$
Recall that
$\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}=L_{n,i}[Y_{\ell}]-\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}L_{n,i}[X_{\ell j}]\quad\text{and}\quad\beta_{0,j}^{*,\mathrm{uni}}=\mathbb{E}[Y_{i}]-\beta_{1,j}^{*,\mathrm{uni}}\mathbb{E}[X_{ij}]$
Then,
$|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|$ $\leq|L_{n,i}[Y_{\ell}]-\mathbb{E}[Y_{i}]|+O_{a}(1)\,|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|+O_{a}(1)\,|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|$
$+O_{a}(1)\,|L_{n,i}[X_{\ell j}]-\mathbb{E}[X_{ij}]|\,|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|$
By the same concentration bounds, we get that for all $t\in(0,1)$
$\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|>t\right)\leq O_{a}(1)\,\exp(-\Omega_{a}(1)\,n^{3/4}t^{2})$
Similarly, we get that for all $t\in(0,1)$
$\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|>t\right)\leq O_{a}(1)\,\exp(-\Omega_{a}(1)\,n^{3/4}t^{2})$
and
$\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|>t\right)\leq O_{a}(1)\,\exp(-\Omega_{a}(1)\,n^{3/4}t^{2})$
$\square$
The next lemma is about the concentration of the univariate coefficients used in uniPairs.
Lemma 5
Assume A1, A2, A9 and A13. Then, there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for all $(j,k)\in\mathcal{P}$, $n\geq 1$, and $t\in(0,K_{1})$, we have :
$\mathbb{P}\!\left(|\widehat{\beta}_{1,jk}^{\mathrm{uni}}-\beta_{1,jk}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/5}t^{2})$
and
$\mathbb{P}\!\left(|\widehat{\beta}_{0,jk}^{\mathrm{uni}}-\beta_{0,jk}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/5}t^{2})$
Also, for all $(j,k)\in\mathcal{P}$, $i\geq 1$, $n\geq i$ and $t\in(0,K_{1})$, we have :
$\mathbb{P}\!\left(|\widehat{\beta}_{1,jk}^{(-i)\mathrm{uni}}-\beta_{1,jk}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/5}t^{2})$
and
$\mathbb{P}\!\left(|\widehat{\beta}_{0,jk}^{(-i)\mathrm{uni}}-\beta_{0,jk}^{*,\mathrm{uni}}|>t\right)\leq K_{2}\,\exp(-K_{1}\,n^{3/5}t^{2})$
Proof.
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates changed according to the bounds established in Lemma 1. $\square$
The next fact is about the KKT conditions of the non-negative lasso in uniPairs-2stage.
Fact 1
[…]
The KKT conditions imply that there exist multipliers $\widetilde{\nu}_{j}\geq 0$ such that
$\widetilde{\nu}_{j}\,\widehat{\theta}_{j}=0\qquad\forall j\in[p]$
$P_{n}[Y_{i}]-\widehat{\theta}_{0}-\sum_{j=1}^{p}\widehat{\theta}_{j}\,P_{n}[\widehat{\eta}_{j}^{(-i)}]=0$
$\frac{\widetilde{\nu}_{j}+\lambda_{1}}{2}=P_{n}[\widehat{\eta}_{j}^{(-i)}\,Y_{i}]-\widehat{\theta}_{0}\,P_{n}[\widehat{\eta}_{j}^{(-i)}]-\sum_{k=1}^{p}\widehat{\theta}_{k}\,P_{n}[\widehat{\eta}_{k}^{(-i)}\widehat{\eta}_{j}^{(-i)}]\qquad\forall j\in[p]$
A similar fact can be established about the KKT conditions of the non-negative lasso in uniPairs. Before stating the next lemma, for all $j\in[p]$, let $\widehat{\beta}_{0,j}^{(-0)\mathrm{uni}}=\widehat{\beta}_{0,j}^{\mathrm{uni}}$ and define
$M_{j,1}=\max_{i\in[n]}|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}|,\quad M_{j,0}=\max_{i\in[n]}|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}|$
and
$D_{j,1}=\max_{0\leq i\leq n}|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{\star,\mathrm{uni}}|,\quad D_{j,0}=\max_{0\leq i\leq n}|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\beta_{0,j}^{\star,\mathrm{uni}}|$
The next lemma controls the event of having at least one false positive in the UniLasso stage of uniPairs-2stage.
Lemma 6
Assume A1, A2, A4, A5 and A7. Then, there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that after the UniLasso stage in uniPairs-2stage, we have :
$\mathbb{P}\!\Big(\bigcup_{j\notin S_{M}^{\prime}}\{\widehat{\beta}_{j}^{s}\neq 0\}\Big)\leq K_{1}\,np\,\exp\!\Big(-K_{2}\,n^{3/5}\lambda_{1}^{2}\Big)$
Proof.
Fix $j\in[p]$ and assume $\widehat{\beta}_{j}^{s}\neq 0$. Since $\widehat{\beta}_{j}^{s}=\widehat{\theta}_{j}\widehat{\beta}_{1,j}^{\mathrm{uni}}$ and $\widehat{\theta}_{j}\geq 0$, we get $\widehat{\theta}_{j}>0$. So $\widetilde{\nu}_{j}=0$ and we have
$\frac{\lambda_{1}}{2}=P_{n}[\widehat{\eta}_{j}^{(-i)}Y_{i}]-\widehat{\theta}_{0}P_{n}[\widehat{\eta}_{j}^{(-i)}]-\sum_{k=1}^{p}\widehat{\theta}_{k}P_{n}[\widehat{\eta}_{k}^{(-i)}\widehat{\eta}_{j}^{(-i)}]$
Since $P_{n}\!\big[Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\big]=0$, we get
$\frac{\lambda_{1}}{2}=P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j})\big(Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\big)\Big]$
By Cauchy–Schwarz,
$\Big|\frac{\lambda_{1}}{2}\Big|\leq\Big(P_{n}[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})^{2}]\Big)^{1/2}\Big(P_{n}\!\big[(Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)})^{2}\big]\Big)^{1/2}$
[…]
$P_{n}[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})^{2}]$ $\leq 2P_{n}[X_{ij}^{2}]\,\max_{i\in[n]}(\widehat{\beta}_{1,j}^{(-i),\mathrm{uni}})^{2}+4\max_{i\in[n]}(\widehat{\beta}_{0,j}^{(-i),\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}})^{2}+4(\beta_{0,j}^{*,\mathrm{uni}}-\widehat{\beta}_{0,j}^{\mathrm{uni}})^{2}$
$\leq 2P_{n}[X_{ij}^{2}]M_{j,1}^{2}+8D_{j,0}^{2}$
Thus,
$\Big|\frac{\lambda_{1}}{2}\Big|\leq(P_{n}[Y_{i}^{2}])^{1/2}\Big(2P_{n}[X_{ij}^{2}]M_{j,1}^{2}+8D_{j,0}^{2}\Big)^{1/2}$
Hence,
$\mathbb{P}(\widehat{\theta}_{j}>0)$ $\leq\mathbb{P}\!\left(\frac{\lambda_{1}^{2}}{4}\leq P_{n}[Y_{i}^{2}]\,\big(2P_{n}[X_{ij}^{2}]M_{j,1}^{2}+8D_{j,0}^{2}\big)\right)$
$\leq\mathbb{P}\!\left(P_{n}[Y_{i}^{2}]\,\geq 2\mathbb{E}[Y_{i}^{2}]\right)+\mathbb{P}\!\left(P_{n}[X_{ij}^{2}]\,\geq 2\mathbb{E}[X_{ij}^{2}]\right)+\mathbb{P}\!\left(\frac{\lambda_{1}^{2}}{4}\leq 2\mathbb{E}[Y_{i}^{2}]\,\big(4\mathbb{E}[X_{ij}^{2}]M_{j,1}^{2}+8D_{j,0}^{2}\big)\right)$
$\leq\mathbb{P}\!\left(P_{n}[Y_{i}^{2}]\,\geq 2\mathbb{E}[Y_{i}^{2}]\right)+\mathbb{P}\!\left(P_{n}[X_{ij}^{2}]\,\geq 2\mathbb{E}[X_{ij}^{2}]\right)+\mathbb{P}\!\left(M_{j,1}^{2}\geq\frac{\lambda_{1}^{2}}{64\mathbb{E}[Y_{i}^{2}]\mathbb{E}[X_{ij}^{2}]}\right)$
$+\mathbb{P}\!\left(D_{j,0}^{2}\geq\frac{\lambda_{1}^{2}}{128\mathbb{E}[Y_{i}^{2}]}\right)$
The first two terms are upper bounded respectively by $O_{a}(1)\,\exp(-\Omega_{a}(1)n^{3/5})$ and $O_{a}(1)\,\exp(-\Omega_{a}(1)n)$. By A5, if $j\notin S_{M}$, then $\Omega_{a}(1)|\beta_{1,j}^{*,\mathrm{uni}}|\leq\lambda_{1}\leq O_{a}(1)$, so by Lemma  4
$\mathbb{P}\!\left(M_{j,1}^{2}\geq\frac{\lambda_{1}^{2}}{64\mathbb{E}[Y_{i}^{2}]\mathbb{E}[X_{ij}^{2}]}\right)$ $\leq n\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i),\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
$\leq n\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i),\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
$\leq O_{a}(1)n\,\exp\!\big(-\Omega_{a}(1)\,n^{3/4}\lambda_{1}^{2}\big)$
Similarly,
$\mathbb{P}\!\left(D_{j,0}^{2}\geq\frac{\lambda_{1}^{2}}{128\mathbb{E}[Y_{i}^{2}]}\right)$ $\leq n\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{(-i),\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)+\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
$\leq O_{a}(1)n\,\exp\!\big(-\Omega_{a}(1)\,n^{3/4}\lambda_{1}^{2}\big)$
Therefore, we get that
$\mathbb{P}\!\Big(\bigcup_{j\notin S_{M}^{\prime}}\{\widehat{\beta}_{j}^{s}\neq 0\}\Big)\leq\mathbb{P}\!\Big(\bigcup_{j\notin S_{M}^{\prime}}\{\widehat{\theta}_{j}\neq 0\}\Big)\leq O_{a}(1)\,np\,\exp\!\Big(-O_{a}(1)\,n^{3/5}\lambda_{1}^{2}\Big)$
$\square$
The next lemma controls the event of having at least one false positive in the UniLasso stage of uniPairs.
Lemma 7
Assume A1, A2, A4, A9, A10, A11 and A13. Then, there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that after the UniLasso stage in uniPairs, we have :
$\mathbb{P}\!\Big(\Big(\bigcup_{j\notin S_{M}^{\prime}}\{\widehat{\beta}_{j}^{s}\neq 0\}\Big)\bigcup\Big(\bigcup_{(j,k)\notin S_{I}}\{\widehat{\beta}_{j,k}^{s}\neq 0\}\Big)\Big)\leq K_{1}\,np^{2}\,\exp\!\Big(-K_{2}\,n^{3/5}\lambda_{1}^{2}\Big)$
Proof.
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates changed according to the bounds established in Lemma 5. $\square$
Before stating our next lemma, we introduce the following notation. Let
$\Delta=\max\big(\max_{k\in S_{M}}|\beta_{k}^{*}-\widehat{\beta}_{k}^{s}|\,,\,|\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s}|\big),\quad\widetilde{Q_{1}}=\max_{k\in S_{M}\cup\{0\}}|P_{n}[\varepsilon_{i}^{\prime}X_{ik}]|$
and
$Q_{1}=\max_{k\in S_{M}^{\prime}}P_{n}[|\varepsilon_{i}^{\prime}X_{ik}|],\qquad Q_{2}=\max_{k\in S_{M}^{\prime}}P_{n}[X_{ik}^{2}],\qquad Q_{3}=\max_{k\in S_{M}^{\prime}}P_{n}[|X_{ik}|]$
where $X_{i,0}=1$. Also, let $U_{1}=\min_{j\in S_{M}}|\beta_{1,j}^{*,\mathrm{uni}}|$ and define
$D_{1}=\max_{j\in S_{M}}D_{j,1},\quad M_{1}=\max_{j\in S_{M}}M_{j,1},\quad D_{0}=\max_{j\in S_{M}}D_{j,0},\quad M_{0}=\max_{j\in S_{M}}M_{j,0}$
The next lemma concludes the proof of Theorem 1. Let $E=\bigcup_{j\notin S_{M}^{\prime}}\{\widehat{\theta}_{j}\neq 0\}$ and $\delta=\max_{k\in S_{M}^{\prime}}|\mathbb{E}(\epsilon_{i}^{\prime}X_{ik})|$.
Lemma 8
Assume A1, A2, A3, A4, A5, and A7. Then, there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for all $n$ large enough,
$\mathbb{P}(E^{c}\cap(\Delta>K_{1}(\lambda_{1}+\delta))\leq K_{1}|S_{M}|n\exp({-K_{2}n^{3/5}}))$
Proof.
Assume $E^{c}$ happens i.e $\forall j\notin S_{M}^{\prime},\widehat{\theta}_{j}=0.$ Let $j\in S_{M}$. We have
$\lambda_{1}\geq 2P_{n}\!\Big[\widehat{\eta}_{j}^{(-i)}\Big(Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\Big)\Big]\qquad\text{and}\qquad\frac{\beta_{j}^{*}}{\beta_{1,j}^{*,\mathrm{uni}}}>0$
Hence by KKT conditions in Fact 1,
$\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}\Bigg(\lambda_{1}-2P_{n}\!\Big[\widehat{\eta}_{j}^{(-i)}\Big(Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\Big)\Big]\Bigg)\geq 0$
i.e
$\frac{\lambda_{1}(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})}{2\beta_{1,j}^{*,\mathrm{uni}}}\geq\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[\widehat{\eta}_{j}^{(-i)}\Big(Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\Big)\Big]$
But,
$P_{n}\!\Big[Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\Big]=0$
So
$\frac{\lambda_{1}(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})}{2\beta_{1,j}^{*,\mathrm{uni}}}\geq\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\Big(Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\Big)\Big]$
Hence,
$\frac{\lambda_{1}(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})}{2\beta_{1,j}^{*,\mathrm{uni}}}$ $\geq\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})(\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s}-\sum_{k\in S_{M}}(\widehat{\beta}_{k}^{s}-\beta_{k}^{*})X_{ik})\Big]$
$\quad+\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\varepsilon_{i}^{\prime}\Big]$
$\quad+\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\Big(\sum_{k\in S_{M}}\widehat{\theta}_{k}\big(\widehat{\beta}_{0,k}^{\mathrm{uni}}-\widehat{\beta}_{0,k}^{(-i)}+(\widehat{\beta}_{1,k}^{\mathrm{uni}}-\widehat{\beta}_{1,k}^{(-i)})X_{ik}\big)\Big)\Big]$
For the second term, we have
$P_{n}\!\Big[\frac{(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\varepsilon_{i}^{\prime}}{\beta_{1,j}^{*,\mathrm{uni}}}\Big]$ $=P_{n}\!\Big[\frac{(\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\varepsilon_{i}^{\prime}}{\beta_{1,j}^{*,\mathrm{uni}}}\Big]+P_{n}\!\Big[\frac{\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}\varepsilon_{i}^{\prime}X_{ij}}{\beta_{1,j}^{*,\mathrm{uni}}}\Big]$
$\geq-2D_{j,0}\frac{P_{n}[|\varepsilon_{i}^{\prime}|]}{|\beta_{1,j}^{*,\mathrm{uni}}|}+P_{n}[\varepsilon_{i}^{\prime}X_{ij}]-D_{j,1}\frac{P_{n}[|\varepsilon_{i}^{\prime}X_{ij}|]}{|\beta_{1,j}^{*,\mathrm{uni}}|}$
So,
$\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\varepsilon_{i}^{\prime}\Big]$ $\geq-2D_{j,0}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}|]+(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})P_{n}[\varepsilon_{i}^{\prime}X_{ij}]-D_{j,1}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}X_{ij}|]$
For the first term, we have
$\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})(\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s}-\sum_{k\in S_{M}}(\widehat{\beta}_{k}^{s}-\beta_{k}^{*})X_{ik})\Big]$
$\geq(\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s})\frac{(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\Big[\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}}\Big]-(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})P_{n}\!\Big[\frac{\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}X_{ij}}{\beta_{1,j}^{*,\mathrm{uni}}}\sum_{k\in S_{M}}(\widehat{\beta}_{k}^{s}-\beta_{k}^{*})X_{ik}\Big]$
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}\!\Big[|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\sum_{k\in S_{M}}|\widehat{\beta}_{k}^{s}-\beta_{k}^{*}||X_{ik}|\Big]$
$\geq-(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})P_{n}\!\Big[X_{ij}\sum_{k\in S_{M}}(\widehat{\beta}_{k}^{s}-\beta_{k}^{*})X_{ik}\Big]-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}\Big[|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}||X_{ij}|\sum_{k\in S_{M}}|\widehat{\beta}_{k}^{s}-\beta_{k}^{*}||X_{ik}|\Big]$
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}|\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s}|(2D_{j,0}+D_{j,1}P_{n}[|X_{ij}|])-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}\Big[|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\sum_{k\in S_{M}}|\widehat{\beta}_{k}^{s}-\beta_{k}^{*}||X_{ik}|\Big]$
$\quad+(\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s})(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})P_{n}\Big[X_{ij}\Big]$
Therefore, for all $j\in S_{M}$,
$\frac{\lambda_{1}(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})}{2\beta_{1,j}^{*,\mathrm{uni}}}-(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})P_{n}\Big[\varepsilon_{i}^{\prime}X_{ij}\Big]+(\beta_{j}^{*}-\widehat{\beta}_{j}^{s})(\widehat{\beta}_{0}^{s}-\beta_{0,M}^{*})P_{n}\Big[X_{ij}\Big]-\sum_{k\in S_{M}}P_{n}\Big[(\widehat{\beta}_{j}^{s}-\beta_{j}^{*})(\widehat{\beta}_{k}^{s}-\beta_{k}^{*})X_{ik}X_{ij}\Big]\geq R_{j}$
where
$R_{j}$ $=-2D_{j,0}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}|]-D_{j,1}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}X_{ij}|]$
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\sum_{k\in S_{M}}\widehat{\theta}_{k}P_{n}\!\Big[|\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\big(|\widehat{\beta}_{0,k}^{\mathrm{uni}}-\widehat{\beta}_{0,k}^{(-i)}|+|\widehat{\beta}_{1,k}^{\mathrm{uni}}-\widehat{\beta}_{1,k}^{(-i)}||X_{ik}|\big)\Big]$
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}\Big[|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}||X_{ij}|\sum_{k\in S_{M}}|\widehat{\beta}_{k}^{s}-\beta_{k}^{*}||X_{ik}|\Big]$
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}|\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s}|(2D_{j,0}+D_{j,1}P_{n}[|X_{ij}|])-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}\Big[|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\sum_{k\in S_{M}}|\widehat{\beta}_{k}^{s}-\beta_{k}^{*}||X_{ik}|\Big]$
So,
$R_{j}$ $\geq-2D_{j,0}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}|]-D_{j,1}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}X_{ij}|]$
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\Biggl\{\sum_{k\in S_{M}}\widehat{\theta}_{k}P_{n}\!\Big[|\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\big(|\widehat{\beta}_{0,k}^{\mathrm{uni}}-\widehat{\beta}_{0,k}^{(-i)}|+|\widehat{\beta}_{1,k}^{\mathrm{uni}}-\widehat{\beta}_{1,k}^{(-i)}||X_{ik}|\big)\Big]$
$\quad+P_{n}\Big[|\widehat{\beta}_{1,j}^{(-i)\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}||X_{ij}|\sum_{k\in S_{M}}|\widehat{\beta}_{k}^{s}-\beta_{k}^{*}||X_{ik}|\Big]+|\beta_{0,M}^{*}-\widehat{\beta}_{0}^{s}|\left(2D_{j,0}+D_{j,1}P_{n}[|X_{ij}|]\right)$
$\quad+P_{n}\Big[|\widehat{\beta}_{0,j}^{(-i)\mathrm{uni}}-\widehat{\beta}_{0,j}^{\mathrm{uni}}|\sum_{k\in S_{M}}|\widehat{\beta}_{k}^{s}-\beta_{k}^{*}||X_{ik}|\Big]\Biggr\}$
So,
$R_{j}$ $\geq-2D_{j,0}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}|]-D_{j,1}\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}P_{n}[|\varepsilon_{i}^{\prime}X_{ij}|]$
$\quad-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}^{s}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\Biggl\{\sum_{k\in S_{M}}\widehat{\theta}_{k}P_{n}\!\Big[(2D_{j,0}+M_{j,1}|X_{ij}|)\big(2D_{k,0}+2D_{k,1}|X_{ik}|\big)\Big]$
$\quad+D_{j,1}\Delta P_{n}\Big[\sum_{k\in S_{M}}|X_{ij}||X_{ik}|\Big]+\Delta(2D_{j,0}+D_{j,1}P_{n}[|X_{ij}|])$
$\quad+2D_{j,0}\Delta P_{n}\Big[\sum_{k\in S_{M}}|X_{ik}|\Big]\Biggr\}$
But, $P_{n}[|X_{ik}X_{ij}|]\leq(P_{n}[X_{ik}^{2}]P_{n}[X_{ij}^{2}])^{\frac{1}{2}}\leq Q_{2}$, so
$R_{j}$ $\geq-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\Big(2D_{j,0}P_{n}[|\varepsilon_{i}^{\prime}|]+D_{j,1}P_{n}[|\varepsilon_{i}^{\prime}X_{ij}|]+4D_{j,0}\Theta_{0}+4D_{j,0}Q_{3}\Theta_{1}+2M_{j,1}Q_{3}\Theta_{0}+2M_{j,1}Q_{2}\Theta_{1}$
$+\,D_{j,1}\Delta\,|S_{M}|\,Q_{2}+\,\Delta\,(2D_{j,0}+D_{j,1}Q_{3})+\,2D_{j,0}\Delta\,|S_{M}|\,Q_{3}\Big)$
where $\Theta_{1}=\sum_{k\in S_{M}}\widehat{\theta}_{k}D_{k,1}$ and $\Theta_{0}=\sum_{k\in S_{M}}\widehat{\theta}_{k}D_{k,0}$. Therefore
$R_{j}$ $\geq-\frac{|\beta_{j}^{*}-\widehat{\beta}_{j}|}{|\beta_{1,j}^{*,\mathrm{uni}}|}\Big(2D_{j,0}\,Q_{1}+D_{j,1}\,Q_{1}+4D_{j,0}\Theta_{0}+4D_{j,0}Q_{3}\Theta_{1}+2M_{j,1}Q_{3}\Theta_{0}+2M_{j,1}Q_{2}\Theta_{1}$
$+\,D_{j,1}\Delta\,|S_{M}|\,Q_{2}+\,\Delta\,(2D_{j,0}+D_{j,1}Q_{3})+\,2D_{j,0}\Delta\,|S_{M}|\,Q_{3}\Big)$
For $j\in S_{M}$, let $\theta_{j}^{*}=\beta_{j}^{*}/\beta_{1,j}^{*,\mathrm{uni}}$. Then,
$|\widehat{\theta}_{j}-\theta_{j}^{*}|$ $\leq\frac{|\widehat{\beta}_{j}-\beta_{j}^{*}|}{|\widehat{\beta}_{1,j}|}+|\beta_{j}^{*}|\big|\frac{1}{\beta_{1,j}^{*,\mathrm{uni}}}-\frac{1}{\widehat{\beta}_{1,j}}\big|$
We have
$\mathbb{P}\!\left(2D_{1}>U_{1}\right)\leq O_{a}(1)|S_{M}|n\,\exp(-\Omega_{a}(1)\,n^{3/4})$
And if $2D_{1}\leq U_{1}$, then
$|\widehat{\theta}_{j}-\theta_{j}^{*}|$ $\leq\frac{2\Delta}{|\beta_{1,j}^{*,\mathrm{uni}}|}+|\beta_{j}^{*}|\frac{2D_{j,1}}{(\beta_{1,j}^{*,\mathrm{uni}})^{2}}$
By A4 and A7, we have $\max_{j\in S_{M}}|\theta_{j}^{*}|=O_{a}(1)$, and so
$\sum_{j\in S_{M}}R_{j}$ $\geq-\frac{O_{a}(1)\Delta}{U_{1}}\Biggl\{(D_{0}+D_{1})Q_{1}+(D_{0}+Q_{3}M_{1})(\frac{\Delta D_{0}}{U_{1}}+\frac{D_{0}D_{1}}{U_{1}^{2}}+D_{0})$
$+(D_{0}Q_{3}+M_{1}Q_{2})(\frac{\Delta D_{1}}{U_{1}}+\frac{D_{1}^{2}}{U_{1}^{2}}+D_{1})+D_{1}\Delta Q_{2}+\Delta D_{0}+\Delta D_{1}Q_{3}+D_{0}\Delta Q_{3}\Biggr\}$
Define $\widetilde{\beta_{k}^{*}}=\beta_{k}^{*}$ if $k>0$ else $\widetilde{\beta_{0}^{*}}=\beta_{0,M}^{*}$. Therefore,
$\sum_{j,k\in S_{M}^{\prime}}(\widetilde{\beta_{j}^{*}}-\widehat{\beta}_{j}^{s})(\widetilde{\beta_{k}^{*}}-\widehat{\beta}_{k}^{s})P_{n}[X_{ik}X_{ij}]\leq O_{a}(1)\big(\Delta\lambda_{1}+\Delta\widetilde{Q_{1}}+\Delta A_{1}+\Delta^{2}A_{2}\big)$
where
$A_{1}=\widetilde{D}\widetilde{Q}+\widetilde{D}^{3}+\widetilde{D}^{2}+\widetilde{D}^{2}\widetilde{Q}^{2}+\widetilde{D}^{3}\widetilde{Q}+\widetilde{D}^{2}\widetilde{Q}+\widetilde{D}\widetilde{Q}^{2}\quad\text{and}\quad A_{2}=\widetilde{D}\widetilde{Q}+\widetilde{D}+\widetilde{D}^{2}+\widetilde{D}\widetilde{Q}^{2}+\widetilde{D}^{2}\widetilde{Q}$
with $\widetilde{D}=\max(D_{0},D_{1})$ and $\widetilde{Q}=\max(M_{1},Q_{1},Q_{2},Q_{3})$. Let $A_{3}=(\widetilde{D}+\widetilde{D}^{2}+\widetilde{D}^{3})(\widetilde{Q}+\widetilde{Q}^{2})$. Then,
$\sum_{j,k\in S_{M}^{\prime}}(\widetilde{\beta_{j}^{*}}-\widehat{\beta}_{j}^{s})(\widetilde{\beta_{k}^{*}}-\widehat{\beta}_{k}^{s})P_{n}[X_{ik}X_{ij}]\\ \leq O_{a}(1)\big(\Delta\lambda_{1}+\Delta\widetilde{Q_{1}}+\Delta A_{3}+\Delta^{2}A_{3}\big)$
$\widehat{\eta}_{M}$ is the smallest eigenvalue of the PSD matrix $(P_{n}[X_{ik}X_{ij}])_{j,k\in S_{M}^{\prime}}$, so
$\widehat{\eta}_{M}\Delta^{2}\leq O_{a}(1)\big(\Delta\lambda_{1}+\Delta\widetilde{Q_{1}}+\Delta A_{3}+\Delta^{2}A_{3}\big)$
When $\Delta>0$, we get
$\Delta\leq O_{a}(1)\frac{\lambda_{1}+\widetilde{Q_{1}}+A_{3}}{\widehat{\eta}_{M}-O_{a}(1)A_{3}}$
We have for all $t>0$
$\mathbb{P}(\tilde{D}>t)\leq O_{a}(1)|S_{M}|n\exp(-\Omega_{a}(1)n^{3/4}t^{2})$
For all $A$ larger than $2\max_{k}\mathbb{E}(|\epsilon_{i}^{\prime}X_{ik}|)+2\max_{k}\mathbb{E}(X_{ik}^{2})+2\max_{k}\mathbb{E}(|X_{ik}|)+2\max_{k}|\beta_{1,k}^{*,\text{uni}}|$ which is $O_{a}(1)$, we have
$\mathbb{P}(\tilde{Q}>A)$ $\leq O_{a}(1)(n+1)|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A^{2})+O_{a}(1)(|S_{M}|+1)\exp(-\Omega_{a}(1)nA^{2})$
$+O_{a}(1)(|S_{M}|+1)\exp(-\Omega_{a}(1)n^{3/4}A)$
Hence, $\mathbb{P}(\tilde{Q}>A)\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)$. Recall that $\widetilde{Q_{1}}=\max_{k\in S_{M}^{\prime}}|P_{n}[\varepsilon_{i}^{\prime}X_{ik}]|$ and $\delta=\max_{k\in S_{M}^{\prime}}|\mathbb{E}(\epsilon_{i}^{\prime}X_{ik})|$. So,
$|\widetilde{Q_{1}}-\delta|\leq\max_{k\in S_{M}^{\prime}}|P_{n}[\varepsilon_{i}^{\prime}X_{ik}]-\mathbb{E}(\epsilon_{i}^{\prime}X_{ik})|$
Hence,
$\mathbb{P}\!\left(\big|\widetilde{Q_{1}}-\delta\big|>t\right)\leq(|S_{M}|+1)\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}\right)$
Therefore, since $A_{3}=(\widetilde{D}+\widetilde{D}^{2}+\widetilde{D}^{3})(\widetilde{Q}+\widetilde{Q}^{2})$, we get
$\mathbb{P}(A_{3}>(t+t^{2}+t^{3})(A+A^{2}))$ $\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$
So for all $t\in(0,\Omega_{a}(1))$ and $A=\Omega_{a}(1)$, we have
$\mathbb{P}(A_{3}>tA^{2})$ $\leq O_{a}(1)n|S_{M}|\exp(-\Omega_{a}(1)n^{3/4}A)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$
Therefore, for all $t\in(0,\Omega_{a}(1))$, we have
$\mathbb{P}(A_{3}>t)\leq O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$
Therefore, we conclude that for all $t\in(0,\Omega_{a}(1))$,
$\mathbb{P}\Big(E^{c}\cap\big(\Delta>O_{a}(1)\frac{\lambda_{1}+\delta+t}{\eta^{*}_{M}-O_{a}(1)n^{-\epsilon/2}-O_{a}(1)t}\big)\Big)$
$\leq O_{a}(1)|S_{M}|\exp\!\left(-\Omega_{a}(1)\,t\,n^{3/4}\right)+O_{a}(1)|S_{M}|\exp\!\left(-\Omega_{a}(1)n^{3/4}\right)+O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/4}t^{2}})$
$+O_{a}(1)\exp(-\Omega_{a}(1)n^{1-\epsilon})$
By taking $t=n^{-3/40}=o(1)$ and $\epsilon=1/4$, we get for all $n$ large enough
$\mathbb{P}(E^{c}\cap(\Delta>O_{a}(1)(\lambda_{1}+\delta))\leq O_{a}(1)|S_{M}|n\exp({-\Omega_{a}(1)n^{3/5}}))$
$\square$
Combining Lemma 6 and Lemma 8 and noting that $|S_{M}|=O_{a}(1)$ by A7, we conclude Theorem 1. The next lemma concludes the proof of Theorem 2. Let
$E^{\prime}=\Big(\bigcup_{j\notin S_{M}^{\prime}}\{\widehat{\beta}_{j}^{s}\neq 0\}\Big)\bigcup\Big(\bigcup_{(j,k)\notin S_{I}}\{\widehat{\beta}_{j,k}^{s}\neq 0\}\Big)$
and
$\Delta^{\prime}=\max\big(\max_{k\in S_{M}}|\beta_{k}^{*}-\widehat{\beta}_{k}^{s}|\,,\,|\beta_{0}^{*}-\widehat{\beta}_{0}^{s}|\,,\,\max_{(j,k)\in S_{I}}|\gamma_{jk}^{*}-\widehat{\beta}_{jk}^{s}|\big)$
Lemma 9
Assume A1, A2, A4, A8, A9, A10, A11, and A13. Then, there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for all $n$ large enough,
$\mathbb{P}(E^{\prime c}\cap(\Delta^{\prime}>K_{1}\lambda)\leq K_{1}|S_{A}|n\exp({-K_{2}n^{1/3}}))$
Proof.
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates adapted according to the bounds established in Lemmas 1, 3, 4 and 5. In particular, the following rates are modified :
  • Let $D_{1}^{\prime}=\max\Big(D_{1},\max_{(j,k)\in S_{I},0\leq i\leq n}|\widehat{\beta}_{1,jk}^{(-i)\mathrm{uni}}-\beta_{1,jk}^{\star,\mathrm{uni}}|\Big)$ and similarly define $D_{0}^{\prime}$. Also, let $U_{1}^{\prime}=\min\Big(U_{1},\min_{(j,k)\in S_{I}}|\beta_{1,jk}^{*,\mathrm{uni}}|\Big)$. Then,
    $\mathbb{P}\!\left(2D_{1}^{\prime}>U_{1}^{\prime}\right)\leq O_{a}(1)|S_{A}|n\,\exp(-\Omega_{a}(1)\,n^{3/5})$
  • Let $\widetilde{D^{\prime}}=\max(D_{0}^{\prime},D_{1}^{\prime})$. Then for all $t>0$, we have
    $\mathbb{P}(\widetilde{D^{\prime}}>t)\leq O_{a}(1)|S_{A}|n\exp(-\Omega_{a}(1)n^{3/5}t^{2})$
  • Let $\widetilde{Q^{\prime}}=\max(M_{1}^{\prime},Q_{2}^{\prime},Q_{3}^{\prime})$ where $M_{1}=\max\Big(M_{1},\max_{(j,k)\in S_{I},i\in[n]}|\widehat{\beta}_{1,jk}^{(-i)\mathrm{uni}}|\Big)$,
    $Q_{2}^{\prime}=\max\Big(Q_{2},\max_{(j,k)\in S_{I}}P_{n}[X_{ij}^{2}X_{ik}^{2}]\Big),\quad Q_{3}^{\prime}=\max\Big(Q_{3},\max_{(j,k)\in S_{I}}P_{n}[|X_{ij}X_{ik}|]\Big)$
    Then for $A=\Omega_{a}(1)$, we have $\mathbb{P}(\widetilde{Q^{\prime}}>A)\leq O_{a}(1)n|S_{A}|\exp(-\Omega_{a}(1)n^{3/5}A^{4/5})$
  • $\widetilde{Q_{1}}=0$
  • Let $A_{3}^{\prime}=(\widetilde{D^{\prime}}+\widetilde{D^{\prime}}^{2}+\widetilde{D^{\prime}}^{3})(\widetilde{Q^{\prime}}+\widetilde{Q^{\prime}}^{2})$. Then for all $t\in(0,\Omega_{a}(1))$, we have
    $\mathbb{P}(A_{3}>t)\leq O_{a}(1)|S_{A}|n\exp({-\Omega_{a}(1)n^{3/5}t^{2}})$
$\square$
Appendix B: Generalization to Binomial GLM and Cox survival model
We provide implementations that extend uniPairs and uniPairs-2stage to the Binomial generalized linear model with logit link (logistic regression) and the Cox proportional hazards model. The algorithmic structure remains the same with two important changes. First, we use the approximation Rad and Maleki [9] of the LOO prediction used in UniLasso since no exact formula exists. Second, the t-tests performed in the Triplet-Scan are changed to likelihood ratio tests with unpenalized GLM fitting instead of OLS.
The model is written in terms of the linear predictor
$\eta=X\beta+\text{offset}$
with inverse link $\mu$. Each model provides a fitted linear predictor $\widehat{\eta}\in\mathbb{R}^{n}$, a log-likelihood $\ell(\eta;y)$ (partial likelihood for Cox), together with its gradient and Hessian evaluated at $\widehat{\eta}$:
$\widehat{g}=g(\widehat{\eta})\;=\;\left.\frac{\partial\ell(\eta;y)}{\partial\eta}\right|_{\eta=\widehat{\eta}}\in\mathbb{R}^{n}\quad\text{and}\quad\widehat{H}=H(\widehat{\eta})\;=\;\left.\frac{\partial^{2}\ell(\eta;y)}{\partial\eta\partial\eta^{\top}}\right|_{\eta=\widehat{\eta}}\in\mathbb{R}^{n\times n}$
When fitting Univariate models inside UniLasso both in uniPairs and uniPairs-2stage, for each $j\in[p]$ and each $(j,k)\in\widehat{\Gamma}$ (the second case only happens for uniPairs), we fit the Univariate GLM models
$\eta_{j}=\beta_{0,j}^{\mathrm{uni}}+\beta_{1,j}^{\mathrm{uni}}\widetilde{X}_{j},\quad\text{and}\quad\eta_{jk}=\beta_{0,jk}^{\mathrm{uni}}+\beta_{1,jk}^{\mathrm{uni}}(\widetilde{X}_{j}\odot\widetilde{X}_{k})$
Exact LOO formulas are unavailable for GLMs, so we use the approximation Rad and Maleki [9]:
$\widehat{\eta}^{(-i)}\;\approx\;\widehat{\eta}_{i}-\frac{\widehat{g}_{i}}{\widehat{H}_{i}\,(1-\widehat{h}_{i})}$
where
$\widehat{h}_{i}=\frac{\widehat{H}_{i}\widetilde{X}_{ij}^{2}}{\sum_{k=1}^{n}\widehat{H}_{i}\widetilde{X}_{kj}^{2}}$
For each pair $(j,k)\in\mathcal{P}$, TripletScan fits the unpenalized GLM
$\eta_{jk}=\beta_{0,jk}+\beta_{j,jk}\widetilde{X}_{j}+\beta_{k,jk}\widetilde{X}_{k}+\beta_{jk,jk}(\widetilde{X}_{j}\odot\widetilde{X}_{k})$
where the likelihood corresponds to the Binomial or Cox model. To test the interaction coefficient $\beta_{jk,jk}$, we use a likelihood ratio test:
$\Lambda_{jk}=2\big(\ell_{\mathrm{full}}-\ell_{\mathrm{null}}\big)\quad\text{and}\quad p_{jk}=\mathbb{P}\!\big(\chi^{2}_{1}\geq\Lambda_{jk}\big)$
where the null model excludes the interaction term.
Appendix C: Full simulation results
The next four figures complement the ones shown in Section 6. They show the main-effects model size, the interactions model size, Train $R^{2}$, main-effects FDR, interactions FDR, main-effects coverage, and interactions coverage, as well as the Jaccard index between the predicted active sets of uniPairs and uniPairs-2stage.
Figure 8: Main-effects model size for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$. The red dashed line marks the true number of active main effects. Glinternet is plotted against the right y-axis while uniPairs, uniPairs-2stage and Sprinter use the left y-axis.
In Figure 8, we see that uniPairs and uniPairs-2stage produce main-effects model sizes that are close to the true number of active main effects. Their selection tends to be slightly above the truth. Sprinter tends to select more main-effects at high SNR levels but overall stays close to the two variants. In contrast, Glinternet selects a very large number of main-effects in every scenario, often exceeding one hundred even when only six main effects are truly active.
Figure 9: Interactions model size for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$. The red dashed line marks the true number of active interactions. Glinternet is plotted against the right y-axis while uniPairs, uniPairs-2stage and Sprinter use the left y-axis.
In Figure 9, we see that uniPairs and uniPairs-2stage consistently select very few interaction terms, typically very close to the true number of active pairs. This holds across all structures and SNR levels and their behavior remains very stable. The selected interaction counts for Sprinter are often several times larger than the truth. Glinternet selects extremely large numbers of interactions, in the order of thousands, irrespective of structure and SNR level. Even in the main-effects only case, where the true number of interactions is zero, Glinternet produces a large interaction set. This explains its high coverage in Figure 14 which is achieved through aggressive over-selection.
Figure 10: Train $R^{2}$ for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 10, we see that Train $R^{2}$ increases with SNR for all methods, as expected. uniPairs and uniPairs-2stage obtain slightly lower Train $R^{2}$ than Glinternet and Sprinter, especially at high SNR levels, but with much sparser models as seen in Figure 4.
Figure 11: Main-effects false discovery rate (FDR) for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 11, we see that Glinternet consistently exhibits the highest main-effects FDR, independent of structure and SNR level. This is consistent with its over-selection of main-effects as seen in Figure 8 and shows that many of its selected main-effects are false positives. uniPairs and uniPairs-2stage maintain substantially lower FDR which is on average slightly lower than that of Sprinter.
Figure 12: Interaction false discovery rate (FDR) for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 12, we see that both Sprinter and Glinternet exhibit high interactions FDR, which is consistent at least for Glinternet with its over-selection of interactions as seen in Figure 9. In contrast, uniPairs and uniPairs-2stage maintain a lower interactions FDR, which tends to decrease slightly as SNR increases. This pattern holds across structures and SNR levels, showing that uniPairs and uniPairs-2stage identify interactions more conservatively, leading to lower FDR and more interpretable models.
Figure 13: Main-effects coverage for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$. Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 13, Glinternet consistently achieves the highest main-effects coverage, but at the expense of many false positives as seen in Figure 8. uniPairs and uniPairs-2stage maintain good coverage, typically around $0.6$ at low SNR, and around $0.7$ at high SNR. Sprinter has considerably lower coverage than the other methods, particulary at low SNR.
Figure 14: Interaction coverage for $(n,p)=(300,400)$ aggregated over $\rho\in\{0,0.2,0.5,0.8,1\}$.Each bar shows mean $\pm$ one standard error across +200 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 14, we see that Glinternet achieves the highest interaction coverage in every setting, but at the cost of extremely high FDR and very large model size as seen in Figures 12 9. uniPairs and uniPairs-2stage achieve moderate interactions coverage, but with far fewer false positives as seen in Figure 12. Sprinter shows the lowest overall interactions coverage.
Figure 15: Jaccard index $\frac{|A\cap B|}{|A\cup B|}$ between the predicted models of uniPairs and uniPairs-2stage for $n=300,p=400$ and $\rho=0.8$. Each bar shows mean $\pm$ one standard error across 50 replicates. Rows correspond to structures and columns to SNR levels $(0.5,1,3)$.
In Figure 15, we see that the Jaccard index between the models selected by uniPairs and uniPairs-2stage is high for both main-effects and interactions. Across all structures and SNR levels, interaction sets exhibit particularly high similarity. For main effects, the agreement is slightly lower than for interactions.